### v2

# Leveraging Multi-facet Paths for Heterogeneous Graph Representation Learning

Abstract.
Recent advancements in heterogeneous GNNs have enabled significant progress in embedding nodes and learning relationships across diverse tasks. However, traditional methods rely heavily on meta-paths grounded in node types, which often fail to encapsulate the full complexity of node interactions, leading to inconsistent performance and elevated computational demands. To address these challenges, we introduce MF2Vec, a novel framework that shifts focus from rigid node-type dependencies to dynamically exploring shared facets across nodes, regardless of type. MF2Vec constructs multi-faceted paths and forms homogeneous networks to learn node embeddings more effectively. Through extensive experiments, we demonstrate that MF2Vec achieves superior performance in node classification, link prediction, and node clustering tasks, surpassing existing baselines. Furthermore, it exhibits reduced performance variability due to meta-path dependencies and achieves faster training convergence. These results highlight its capability to analyze complex networks comprehensively. The implementation of MF2Vec is publicly available at https://anonymous.4open.science/r/MF2Vec-6ABC.
1. Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for capturing complex structures and relationships within graph data, with applications spanning numerous domains (Kipf and Welling, 2016; Hamilton et al., 2017; Velickovic et al., 2017). As a specialized branch of GNNs, Heterogeneous Graph Neural Networks (HGNNs) (Fu and King, 2024; Jin et al., 2022; Guan et al., 2023; Fu et al., 2020; Wang et al., 2019) aim to address the unique challenges of modeling graphs with diverse types of nodes and edges by introducing tailored mechanisms for efficient learning. Among these, relation-based HGNNs enhance learning by aggregating messages along edges categorized by relation types, yet they often encounter limitations due to the deep network structures required to process multi-hop relations, leading to over-smoothing and diminished performance (Fu and King, 2024).
To mitigate these challenges, meta-path-based HGNNs leverage meta-paths to construct homogeneous subgraphs, reducing the need for multi-hop message passing. For instance, the “MAM” meta-path on the IMDB dataset links Movie and Actor nodes, capturing collaborative relationships between movies via shared actors (Fu and King, 2024; Fu et al., 2020; Wang et al., 2021). Despite their efficiency, such meta-path-based approaches are constrained by rigid node-type combinations, which give rise to several critical limitations.
First, meta-paths can only reflect specific relationships between certain types of nodes, which limits their ability to capture a wide range of relationships in a detailed manner. For example, the “MAM” meta-path captures collaborations between movies that share the same actor but fails to represent more detailed relationships beyond it such as sharing genres or countries. Second, as depicted in Figure 1 (left), the performance of meta-path-based models fluctuates depending on the selected meta-path, resulting in instability across datasets and tasks (Sun et al., 2022; Hu et al., 2020). Third, as shown in Figure 1 (right), the exponential increase in meta-path combinations and lengths with diverse intermediate nodes amplifies computational costs and exploration complexity. Even with advanced learnable meta-path approaches like GTN (Yun et al., 2019) and MHNF (Sun et al., 2022), these issues persist due to their reliance on fixed node-type configurations.
This paper introduces a novel Multi-Facet Path approach to address these limitations. The concept of a facet, representing distinct attributes or properties of nodes (e.g., filming location or production year for a movie), is expanded from prior works on homogeneous graphs (Liu et al., 2019; Park et al., 2020; Chen et al., 2021) to encompass shared characteristics across diverse node types in heterogeneous graphs. By shifting from node-type-based paths to facet-based constructions, the proposed method offers a flexible and granular modeling framework that overcomes the rigidity and inefficiency of traditional meta-path approaches.
Through dynamic facet embeddings, which are learned and updated in an end-to-end manner, the Multi-Facet Path approach enables the representation of intricate relationships across nodes, enhancing both efficiency and model stability. First, the use of facets allows for the capture of a broader range of relationships, enriching downstream task representations. Second, this method minimizes performance variability by dynamically constructing paths based on facet information rather than fixed meta-paths. Third, the elimination of manual meta-path definitions and the reduction of computational overhead further bolster scalability.
Figure 2 demonstrates that even when following the same meta-path, nodes can be grouped based on different facets. For example, the paths “New York-L. Dicaprio-The Wolf of Wall Street” and “Interstellar-Anne H.-Intern” may simply be grouped as “Movie-Actor-Movie” under meta-path-based methods. However, with a multi-facet approach, the path “New York-L. Dicaprio-The Wolf of Wall Street” can be considered from the perspective of filming location, while the path “Interstellar-Anne H.-Intern” may reflect similarities in IMDb ratings. This approach allows the model to comprehensively capture latent and rich information within the graph, effectively overcoming the limitations of traditional meta-path-based methods.
In MF2Vec, paths are extracted from the adjacency matrix of a heterogeneous graph through random walks without distinguishing them by predefined meta-path types. The node embeddings of the extracted paths are projected into a shared facet space of dimensions. Gumbel Softmax and an attention mechanism are employed to select the most relevant facets for intermediate nodes, ensuring a discrete yet differentiable selection process. These facets characterize the relationships between endpoint nodes and compute path weights that reflect their importance for node connections. Finally, a homogeneous network is constructed by integrating same-typed nodes and fed into a Graph Convolutional Network (GCN) (Kipf and Welling, 2016), which aggregates multi-facet vectors along the paths to generate latent node representations. Key contributions of this work are summarized as follows:
Dynamic Facet Embedding: A novel framework for constructing flexible paths independent of node types.
Efficiency and Performance Enhancement: Reduction in computational costs and enhanced performance for node classification, node clustering, and link prediction tasks
Stable Performance: Significant improvement in stability across datasets and tasks by mitigating dependence on fixed meta-paths.
2. Related Work
2.1. Heterogeneous Graph Neural Network (HGNN)
Graph neural networks (GNNs) effectively handle graph-structured data by learning low-dimensional node representations (Defferrard et al., 2016; Kipf and Welling, 2016; Hamilton et al., 2017; Velickovic et al., 2017). Building on GNNs, heterogeneous graph neural networks (HGNNs) extend representation learning to heterogeneous graphs, enabling tasks like node classification, link prediction, and clustering. Message-passing methods have become prominent, generating embeddings by aggregating neighbor information.
2.1.1. Relation-based HGNNs
Relation-based HGNNs aggregate neighbor information with type-specific weights, eliminating manual meta-path definitions. RGCN (Chen et al., 2019) applies relation-specific GCNs, while RGAT (Busbridge et al., 2019) integrates self-attention for relational interactions. HetSANN (Hong et al., 2020) and HGT (Hu et al., 2020) use meta-relation-based attention, and Simple-HGN (Lv et al., 2021) combines type-level and node-level attention with L2 normalization. While efficient, deeper layers may cause over-smoothing (Li et al., 2018) and performance degradation.
2.1.2. Metapath-based HGNNs
Metapath-based HGNNs analyze complex networks using predefined metapaths, enabling direct links between homogeneous nodes. Metapath2vec (Dong et al., 2017) uses random walks with a skip-gram model (Mikolov et al., 2013), while HAN (Wang et al., 2019) and MAGNN (Fu et al., 2020) aggregate features along metapaths. GTN (Yun et al., 2019) dynamically learns metapaths, and GraphHINGE (Jin et al., 2022) efficiently aggregates interactions. Models like HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MHNF (Sun et al., 2022), and MECCH (Fu and King, 2024) enhance robustness and reduce redundancy. Despite progress, most approaches rely on single-facet paths, limiting their granularity. To address this, MF2Vec introduces a multi-faceted, fine-grained approach for richer node interactions and network understanding.
2.2. Multiple-Vector Network Embedding
Recent studies in graph representation learning focus on learning multiple embeddings for nodes, capturing their multi-faceted nature. PolyDW (Liu et al., 2019) uses matrix factorization-based clustering to assign facet distributions, independently sampling facets for target and context nodes. Splitter (Epasto and Perozzi, 2019) generates multi-facet node embeddings aligned with the original representation. Asp2vec (Park et al., 2020) introduces a differentiable facet selection module and regularization to explore interactions across facets. For multi-relational graphs, r-GAT (Chen et al., 2021) aggregates relation-specific neighborhood information and applies query-aware attention to select relevant facets. However, existing methods are limited to single-node facets in homogeneous graphs. We extend these approaches to heterogeneous graphs by transforming meta-path schemas into diverse facets, linking terminal node facets for richer interaction understanding.
3. Preliminary
This section provides definitions of terminologies used in the current study. The notations and their explanations are summarized in Table 1.
Definition 1 (Heterogeneous Graph). A heterogeneous graph is characterized by a node type mapping function and an edge type mapping function , where and are sets of node and edge types, respectively, with .
Definition 2 (Metapath). A metapath is a sequence (or ), representing a composite relation between node types and .
Definition 3 (Metapath Schema). A metapath schema is a sequence of node types connected by edge types . This schema represents the structural pattern of a metapath and illustrates how different node types are related through specific types of relationships defined by the nodes and edges.
Definition 4 (Multi-facet Node Representation Learning). For a graph with nodes , the goal is to learn a multi-facet node embedding for each node . Here, is the number of facets, and each facet embedding aims to: 1) preserve network structure, 2) capture various facets of , and 3) represent facet interactions.
Definition 5 (Multi-facet Path). Unlike predefined metapaths, a multi-facet path is extracted without manual specification. Intermediate nodes are projected onto multi-facet vectors , forming a multi-facet path . This path integrates information from intermediate nodes through multiple facets, updating node representations based on various latent information beyond node types.
4. Methodology
In this section, we describe our approach, as shown in Figure 3. We begin by extracting paths between homogeneous end nodes using random walks, similar to traditional metapath methods (Dong et al., 2017; Fu et al., 2017). We then project the features of intermediate nodes into multiple facets and aggregate these facets to create multi-facet embedding representations. These representations are used as edge features between path-guided neighbors and the target node. The model pseudocode is provided in Algorithm 1, and the notations are summarized in Table 1.
Notation | Explanation |
-dimensional Euclidean space | |
A heterogeneous graph with a node set and an edge set | |
A subgraph with a certain node type set and a certain node type edge set | |
A node | |
The embedding of node | |
-th multi-facet embedding of node | |
Multi-facet embedding of node | |
Final embedding of node after -th layer | |
The set of whole nodes in the path between node and | |
Intermediate nodes in the path between node and | |
The embedding of the path between node and | |
-th multi-facet embedding of the path between node and | |
Multi-facet embedding of the path between node and | |
Weight of -th facet in the path |
4.1. Extracting Paths Using Random Walk
Given a node type , our method constructs a homogeneous subgraph that identifies direct connections between nodes of type , the same as traditional metapath-based models. Instead of relying on metapath schemas such as APA or APCPA, the method generates a path starting from a node of type and applies a random walk, traversing randomly selected adjacent nodes until encountering a type A node. This approach captures relationships within a specified path length limit, effectively mapping connections within the nodes of a type without being constrained by predefined metapath patterns. Details of the Random Walk algorithm are provided in Algorithm 2.
4.2. Transforming Single Embeddings into Multi-Facet
Our model captures inherent information from all node types before generating a facet embedding. Initially, node information is exchanged through a GNN warm-up phase, which facilitates the aggregation of neighborhood information and enriches the initial node embeddings with structural and semantic context. The node embedding, , is defined during this warm-up phase and serves as the foundation for training the main model. The resulting node embedding is then refined into a multi-dimensional embedding to capture diverse information, as detailed in Equation 1.
Here, represents the multi-facet embedding of node within the -th facet, where is the corresponding weight matrix with dimensions for the -th facet, and represents the total number of facets. The individual facet embeddings are subsequently concatenated into a unified multi-facet embedding, , providing a comprehensive representation.
4.3. Multi-facet Feature embedding Aggregation
To construct a multi-facet feature embedding for paths consisting of multiple nodes, we employ an aggregation function that synthesizes information from the nodes along the path. The mathematical formulation for this aggregation is given by Equation 2.
In the given equation, represents the aggregated multi-facet feature embedding for a connection between nodes and . The set consists of intermediate nodes between and , and is an intermediate node within the path. The function signifies an aggregation operation using mean pooling.
4.4. Discriminative Facet Consideration using Gumbel Softmax
We utilize the Gumbel Softmax (Jang et al., 2016) function to introduce randomness and uncertainty into facet selection, which helps the model explore diverse facet influences in complex relationships. The temperature parameter allows us to control the balance between exploration and exploitation. This approach enables us to capture nuanced facet interactions and dependencies, making our model more adaptable and robust for multi-facet data. Even if the random walk algorithm extracts the paths that are little related to the end nodes, we mitigate their impact by assigning higher weights to the most crucial components using Gumbel softmax. As a result, we effectively operate within constrained environments, even without extracting meaningful paths.
In the given equation, represents Gumbel Softmax. is the final embedding for that combines these probabilities.and is a matrix with dimensions .
4.5. Subgraph Construction
Path-based HGNNs transform the input graph into a new homogeneous subgraph for each node and path, where varies by model. For example, in HAN, represents the metapath--guided neighborhood, while MAGNN defines it as instances of metapath- incident to the node. MECCH constructs subgraphs using most metapaths as illustrated in Figure 4.
Our approach, however, avoids the need for multiple metapaths by leveraging multi-facet information from a single path between nodes. This results in a subgraph that provides richer information than traditional methods. The MF2Vec model efficiently learns complex node relationships by constructing subgraphs as described in Algorithm 3 incorporating multi-dimensional edge features which are multi-facet embedding for a more comprehensive node representation.
4.6. Message Passing
To extract node embeddings, we utilize GCN. In this process, we consider interactions between nodes and utilize path facets as edge features. The formula for calculating graph node embeddings is as follows:
In this equation, represents the activation function, which is ELU, and BN represents the batch normalization. The denotes the path facet feature between node and its neighbor , which is utilized as an edge feature.
4.7. Training Losses
We train our model for two tasks: node classification and link prediction. For node classification, the cross-entropy loss is computed over labeled nodes , measuring the disparity between true labels and predictions :
For link prediction, where node labels are unavailable, we minimize the contrastive loss using positive and negative node pairs:
Here, is the sigmoid function, ensuring effective learning for link prediction.
5. Experiment
In this part, we assess the performance and practicality of our introduced method through tests on link prediction, node classification, and node clustering tasks across five diverse graph datasets. The primary goals of these experiments are to address the following research questions.
RQ1. How well does MF2Vec perform on downstream tasks?
RQ2. How stable is the performance of MF2Vec in downstream tasks?
RQ3. How time-efficient is MF2Vec?
5.1. Experimental Settings
5.1.1. Datasets
We evaluate our model on six heterogeneous graph datasets: DBLP, ACM, IMDB, and Freebase for node classification and clustering, and Yelp and MovieLens for link prediction. Experiments include 1000 trials per node with a walk length of 5, covering typical metapaths (e.g., APA, APCPA) using consistent data splits. Dataset details are summarized in Table 2.
Dataset | # Nodes (# Types) | # Edges | Target (# Labels) | Task |
Yelp | 31,081 (3) | 411,263 | User-Business | LP |
Movielens | 2,672 (4) | 234,695 | User-Movie | LP |
DBLP | 26,128 (4) | 119,783 | Author (4) | NC, CL |
ACM | 11,246 (3) | 17,426 | Paper (3) | NC, CL |
IMDB | 11,519 (3) | 17,009 | Movie (3) | NC, CL |
Freebase | 43,854 (4) | 75,517 | Movie (3) | NC, CL |
5.1.2. Baselines
We compare MF2Vec against state-of-the-art heterogeneous graph neural networks (HGNNs), categorized as follows: (a) Relation-Based HGNNs, including RGCN, RGAT, HGT, HetSANN, and Simple-HGN; and (b) Metapath-Based HGNNs, comprising MP2Vec, HAN, MAGNN, HeCo, MHNF, HGMAE HMSG, and MECCH.
5.1.3. Graph Construction Settings
Following Algorithm 2, we ensured that the node types matched the target node and that the path length was at most 5. For each target node, we performed up to 1000 trials. The number of negative samples was set to 1.
5.1.4. Experimental Settings
To ensure a fair comparison among all heterogeneous graph neural networks (HGNNs), the node embedding dimension was fixed to 64. The learning rate was set to , weight decay to , and the dropout rate to 0.5. For each baseline model, hyperparameters such as batch size, attention heads, dropout rate, learning rate, and epochs—excluding the node embedding dimension—followed the default settings specified in the original papers. Training was conducted using the Adam optimizer with early stopping (patience of 20 epochs) to prevent overfitting. Specifically, for MF2Vec, the temperature () was set to 0.5, and the number of facets () was set to 5. To demonstrate generalization, the experiments were conducted using five seed values (1, 10, 100, 1000, and 10000), with the data split into train (80%), validation (10%), and test (10%). The model seeds were fixed during the experiments, and the average and standard deviation of the results were calculated for all seeds.
5.2. Downstream Task (RQ1)
Dataset | DBLP | ACM | IMDB | Freebase | |||||||||
Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
Relation-based | RGCN | 0.8580.004 | 0.8420.005 | 0.8370.005 | 0.8100.012 | 0.7690.007 | 0.7620.011 | 0.5330.024 | 0.4310.027 | 0.3370.043 | 0.5740.049 | 0.5080.057 | 0.4080.078 |
RGAT | 0.8490.003 | 0.8290.004 | 0.8230.004 | 0.8030.021 | 0.7610.023 | 0.7490.025 | 0.5580.011 | 0.4240.015 | 0.4090.016 | 0.6220.027 | 0.5530.033 | 0.4870.037 | |
HGT | 0.9260.017 | 0.8890.026 | 0.8990.022 | 0.7790.016 | 0.7410.016 | 0.7090.024 | 0.5400.019 | 0.3960.024 | 0.3860.022 | 0.5780.030 | 0.4820.034 | 0.4270.037 | |
HetSANN | 0.9500.006 | 0.9310.007 | 0.9250.008 | 0.7800.023 | 0.7360.025 | 0.7110.027 | 0.6130.006 | 0.4870.010 | 0.4840.007 | 0.6460.021 | 0.5750.027 | 0.5210.030 | |
Simple-HGN | 0.9470.011 | 0.9280.011 | 0.9200.015 | 0.8400.037 | 0.8080.033 | 0.7930.045 | 0.6120.035 | 0.4970.038 | 0.4810.045 | 0.6710.014 | 0.6300.017 | 0.5480.031 | |
Metapath-based | MP2Vec | 0.5860.020 | 0.3990.025 | 0.3720.026 | 0.7310.010 | 0.6630.010 | 0.6330.014 | 0.5110.011 | 0.3970.016 | 0.3780.013 | 0.6050.019 | 0.6030.022 | 0.5060.030 |
HAN | 0.9370.003 | 0.9190.003 | 0.9090.004 | 0.7850.011 | 0.7620.009 | 0.7290.014 | 0.5500.020 | 0.4260.021 | 0.3840.034 | 0.6650.011 | 0.6430.014 | 0.5170.033 | |
MAGNN | 0.9720.003 | 0.9260.003 | 0.9150.003 | 0.9450.014 | 0.8060.031 | 0.7770.049 | 0.7390.013 | 0.5340.032 | 0.4920.070 | 0.7440.015 | 0.6740.019 | 0.5180.063 | |
HeCo | 0.9780.003 | 0.8940.013 | 0.8860.014 | 0.8310.019 | 0.7450.012 | 0.7030.013 | 0.5440.006 | 0.3850.014 | 0.3600.004 | 0.7530.021 | 0.6520.021 | 0.5150.033 | |
MHNF | 0.9250.004 | 0.8960.006 | 0.8860.006 | 0.8830.009 | 0.7600.012 | 0.7110.016 | 0.5300.042 | 0.4080.042 | 0.3240.071 | 0.5120.013 | 0.4510.014 | 0.2650.053 | |
HGMAE | 0.9480.018 | 0.8740.121 | 0.8290.084 | 0.8120.011 | 0.7320.012 | 0.7360.019 | 0.6140.019 | 0.4270.016 | 0.3530.033 | 0.7270.030 | 0.6220.030 | 0.4780.042 | |
HMSG | 0.9780.003 | 0.9010.020 | 0.8940.021 | 0.9230.013 | 0.7950.019 | 0.7850.020 | 0.6470.021 | 0.4720.017 | 0.4610.022 | 0.7440.026 | 0.6790.040 | 0.5310.038 | |
MECCH | 0.9910.133 | 0.9410.165 | 0.9370.176 | 0.9160.133 | 0.7900.165 | 0.7930.176 | 0.6570.133 | 0.5120.165 | 0.5020.176 | 0.7730.133 | 0.6480.165 | 0.5880.176 | |
MF2Vec | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
5.2.1. Node Classification
In our node classification experiments, we convert nodes into low-dimensional vector embeddings and apply a Softmax function to obtain probability distributions over label categories. We evaluate the model using Micro-F1, Macro-F1, and AUC metrics, providing a balanced assessment of classification performance. These metrics are used to compare our model, MF2Vec, against baseline models on the DBLP, ACM, IMDB and Freebase datasets, with results presented in Table LABEL:table_1.
Our approach consistently outperforms baseline models across all metrics, demonstrating the effectiveness of MF2Vec. The significant performance improvement, particularly on the IMDB dataset, highlights the model’s ability to capture complex relational data through multi-facet vectors. This suggests that MF2Vec effectively reveals nuanced relationships within heterogeneous information networks (HINs), enhancing understanding beyond node types.
5.2.2. Link Prediction
Dataset | Yelp | Movielens | ||||
Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
RGCN | 0.6350.141 | 0.6210.103 | 0.6140.134 | 0.7410.003 | 0.6640.003 | 0.6640.003 |
RGAT | 0.6970.104 | 0.6650.081 | 0.6550.086 | 0.6870.115 | 0.6360.101 | 0.6300.104 |
HGT | 0.8320.004 | 0.7710.004 | 0.7700.005 | 0.7550.136 | 0.7110.116 | 0.7110.115 |
HetSANN | 0.7030.017 | 0.6890.013 | 0.6880.014 | 0.7740.013 | 0.7720.013 | 0.7710.013 |
Simple-HGN | 0.8240.016 | 0.7570.018 | 0.7570.018 | 0.8200.007 | 0.7460.012 | 0.7450.015 |
MP2Vec | 0.7180.003 | 0.6580.002 | 0.6640.003 | 0.7980.004 | 0.6620.004 | 0.6690.002 |
HAN | 0.6270.102 | 0.5280.076 | 0.3990.140 | 0.8220.004 | 0.6420.050 | 0.6040.072 |
MAGNN | 0.8350.005 | 0.7730.002 | 0.7740.003 | 0.8580.010 | 0.7760.008 | 0.7810.009 |
HeCo | 0.8380.003 | 0.7670.004 | 0.7550.002 | 0.8390.003 | 0.7550.004 | 0.7710.005 |
HMSG | 0.8510.003 | 0.7840.002 | 0.8030.004 | 0.8140.006 | 0.7440.006 | 0.7530.013 |
MECCH | 0.8670.012 | 0.7820.013 | 0.7740.015 | 0.7520.006 | 0.6900.015 | 0.6340.040 |
MF2Vec | 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 |
In Movielens and Yelp, link prediction is framed as a binary classification task to determine the presence of an edge in the original graph. During training, validation, and testing, negative edges are created by replacing the destination node in positive edges with a randomly chosen node that is not connected to the source node. The link probability is computed using the dot product of node embeddings: , where is the sigmoid function.
Table 4 presents the AUC, Micro-F1, and Macro-F1 scores for our model and baselines on Yelp and Movielens. MF2Vec performs competitively on Yelp (against MECCH) and outperforms all baselines on MovieLens. These results highlight the advantage of using multi-facet vectors derived from node paths for link prediction, and demonstrate the efficacy of our model in updating embeddings without relying on node types.
Dataset | DBLP | ACM | IMDB | Freebase | ||||
Model | NMI | ARI | NMI | ARI | NMI | ARI | NMI | ARI |
RGCN | 0.0168 | 0.0670 | 0.3612 | 0.2809 | 0.0088 | 0.0056 | 0.0228 | 0.0443 |
RGAT | 0.4435 | 0.4097 | 0.3976 | 0.3140 | 0.0477 | 0.0143 | 0.0327 | 0.0387 |
HGT | 0.7722 | 0.8250 | 0.5064 | 0.5335 | 0.0433 | 0.0459 | 0.0801 | 0.1075 |
HetSANN | 0.752 | 0.7723 | 0.5274 | 0.5263 | 0.0865 | 0.043 | 0.1054 | 0.0967 |
Simple-HGN | 0.7707 | 0.6405 | 0.5309 | 0.5237 | 0.1410 | 0.1180 | 0.1541 | 0.1804 |
MP2Vec | 0.0519 | 0.0330 | 0.1034 | 0.0641 | 0.0065 | 0.0096 | 0.0145 | 0.0286 |
HAN | 0.3533 | 0.1845 | 0.3036 | 0.3528 | 0.0057 | 0.0015 | 0.0211 | 0.0217 |
MAGNN | 0.7964 | 0.8193 | 0.5125 | 0.5317 | 0.1075 | 0.1199 | 0.1872 | 0.1477 |
HeCo | 0.7477 | 0.7872 | 0.5394 | 0.5454 | 0.0049 | 0.0028 | 0.0805 | 0.0793 |
MHNF | 0.7197 | 0.7498 | 0.5385 | 0.5274 | 0.1443 | 0.1467 | 0.1881 | 0.1697 |
HMSG | 0.7278 | 0.7767 | 0.4475 | 0.3630 | 0.0335 | 0.0326 | 0.1257 | 0.1213 |
MECCH | 0.8312 | 0.8739 | 0.4202 | 0.3513 | 0.0070 | 0.0056 | 0.2612 | 0.3127 |
MF2Vec | 0.8606 | 0.8951 | 0.5714 | 0.5790 | 0.1459 | 0.1448 | 0.3083 | 0.3729 |
5.2.3. Node Clustering
We carry out node clustering experiments using the DBLP, ACM, IMDB, and Freebase datasets to assess the embedding quality produced by MF2Vec. In these experiments, similar to node classification, we utilize the embedding vectors of target nodes from the testing set as input for the K-Means model and employ NMI and ARI metrics to gauge performance. As presented in Table 5, our model surpasses the baseline models, except for MHNF (Sun et al., 2022) in ARI on the IMDB dataset, where MF2Vec closely follows it by a small margin, demonstrating our model’s superior ability to generate effective node representations based on multi-facet vectors for node clustering.
5.3. Stability of performance (RQ2)
Dataset | Metric | MAGNN | HMSG | MECCH | MF2Vec |
DBLP | AUC | 0.144 | 0.201 | 0.118 | 0.064 |
Mi-F1 | 0.233 | 0.318 | 0.195 | 0.138 | |
Ma-F1 | 0.251 | 0.277 | 0.197 | 0.139 | |
ACM | AUC | 0.147 | 0.188 | 0.211 | 0.141 |
Mi-F1 | 0.083 | 0.106 | 0.076 | 0.070 | |
Ma-F1 | 0.102 | 0.088 | 0.066 | 0.051 | |
IMDB | AUC | 0.042 | 0.041 | 0.043 | 0.031 |
Mi-F1 | 0.029 | 0.049 | 0.067 | 0.028 | |
Ma-F1 | 0.032 | 0.030 | 0.032 | 0.020 |
This section emphasizes the stability advantage of leveraging multi-facets from paths in our model. Unlike metapath-based models that depend on predefined schemas (e.g., APA, APCPA), MF2Vec achieves consistent performance across different schema types, effectively capturing fine-grained information regardless of metapath type. To validate this, we compare single-metapath performance across all models, including MF2Vec, on the DBLP, ACM, and IMDB datasets.
Table 6 presents the standard deviations in node classification performance using a single metapath. While other models show significant performance variation depending on the chosen metapath, MF2Vec demonstrates remarkable stability. For the DBLP dataset, performance variations highlight the reliance of metapath-based models on predefined paths and the associated time and cost of analysis, which MF2Vec effectively mitigates.
5.4. Time Complexity (RQ3)
Model | Time Complexity |
HetSANN | |
SimpleHGN | |
MAGNN | |
HAN | |
MECCH | |
MF2Vec |
The proposed model’s time complexity is , where is the number of nodes, is the average neighbors, is the number of facets, and is the embedding dimension. Attention adds , and graph convolution contributes , resulting in a total complexity of .
In Table 7, models like MECCH (Fu and King, 2024), HAN (Wang et al., 2019), and MAGNN (Fu et al., 2020) rely on metapaths, incurring higher costs due to growth with metapath length (). MF2Vec avoids this by leveraging facets (), leading to more efficient computation, satisfying .Additionally, we compare MF2Vec with relation-based methods like HetSANN (Hong et al., 2020) and SimpleHGN (Lv et al., 2021), which also avoid metapaths by focusing on edge types.
Figure 5 compares training times (log scale) and Macro F1 scores for MF2Vec and other models on DBLP, IMDB, and Movielens. MAGNN (Fu et al., 2020) and MP2Vec (Dong et al., 2017) are excluded from Movielens due to their excessive training times. MF2Vec achieves the best performance on node classification and link prediction tasks, with significantly lower training times. This efficiency arises from its dynamic use of path characteristics, avoiding the high costs of aggregating multiple paths as seen in MAGNN and MP2Vec, ensuring better scalability and faster convergence without compromising performance.
5.5. Visualization
To evaluate embedding quality, we visualize node representations from various HGNNs on the DBLP dataset using t-SNE (Van der Maaten and Hinton, 2008). Figure 6 shows t-SNE plots for HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MECCH (Fu and King, 2024), and MF2Vec, with colors representing author classes. MF2Vec demonstrates superior clustering within classes and clear decision boundaries, outperforming the more scattered intra-class distributions and weaker inter-class separations of other HGNNs.
6. Study of MF2Vec
To analyze the impact of intrinsic characteristics of MF2Vec with respect to various features including train rati0, the number facets, we additionally conducted extensive experiments.
Dataset | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
Method | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
No. | 0.8130.003 | 0.7430.002 | 0.7520.004 | 0.8320.003 | 0.7540.005 | 0.7430.013 | 0.9740.004 | 0.8680.012 | 0.8540.012 | 0.9120.014 | 0.7880.015 | 0.7590.025 | 0.531 0.022 | 0.378 0.005 | 0.232 0.036 | 0.7310.011 | 0.6460.021 | 0.4750.015 |
Random. | 0.8670.018 | 0.7970.003 | 0.8080.015 | 0.8710.006 | 0.7950.006 | 0.8030.005 | 0.9890.003 | 0.9350.009 | 0.9310.01 | 0.9490.009 | 0.8240.013 | 0.8130.012 | 0.7510.005 | 0.5880.004 | 0.5710.003 | 0.8250.020 | 0.6960.017 | 0.6280.042 |
Gumble. | 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
6.1. Train Ratio ()
As shown in Figure 7, MF2Vec consistently achieves the highest performance across varying training set ratios, demonstrating its robustness and adaptability. While the performance of all models declines as the training ratio decreases, MF2Vec remains its superiority in most cases and retains competitive in the rest. At the lowest ratio of 0.2, it occasionally shows comparable performance to other models. HAN and one baseline were excluded due to significantly lower performance, ensuring clearer visualization. These results highlight MF2Vec’s effectiveness in handling sparse training data.
6.2. Number of Facets ()
Figure 8 shows MF2Vec’s performance across six datasets with varying facet counts. The performance difference relative to “” highlights the advantage of multiple facets. Except for IMDB, optimal results are achieved with five facets, consistently outperforming the single-facet case by effectively leveraging node and path information. Beyond five facets, performance declines, likely due to reduced feature grouping efficiency.
6.3. Role of Facet Selection
As MF2Vec selects the optimal facet for each path in a learnable manner, we conduct an additional experiment to validate its facet selection strategy. Table 8 compares different methods for selecting facet information as edge weights. Our GCN-based results show that the proposed method, using Gumbel-Softmax and attention mechanisms, outperforms all alternatives across datasets. Random facet selection reduces performance, and omitting facets entirely leads to the lowest results. These findings highlight the importance of facet-related information for effective graph learning and the proposed method’s ability to capture complex structures and hidden relationships.
7. Case Study
To evaluate facet balance among node types, we analyzed their distribution as shown in Figure 9. The heatmaps reveal equitable attention weights across node types, with each column representing a specific node and highlighting diverse facet distributions. For example, the Paper node “P1612” emphasizes Facet 1, while the Author node “A3958” distributes attention evenly across facets, reflecting varying significance assigned to facets based on connections. This analysis confirms that nodes either focus on specific facets or distribute attention uniformly, with no strong bias toward any facet, validating the importance of modeling shared facets in heterogeneous graphs.
8. Conclusion
In this study, we introduce MF2Vec, a novel heterogeneous graph neural network model, which generates node representations using multi-facet vectors from various paths, regardless of node types. MF2Vec captures complex latent semantics through these multi-faceted paths, offering a node-type-agnostic embedding approach. Evaluations across six datasets demonstrate that MF2Vec outperforms leading baseline models in both performance and time complexity. Future work could focus on automating path selection within heterogeneous graph neural networks using learnable methods to improve upon current random pre-selection methods, highlighting the potential of dynamic multi-facet path utilization in graph networks.
Appendix
A. Baselines Code and Datasets
MF2Vec is implemented using PyTorch. Table 9 lists the URLs of the authors’ implementations of the compared methods, and Table 10 provides the URLs where the datasets can be downloaded.
B. Baselines
We compare MF2Vec against state-of-the-art relation-based HGNNs and metapath-based HGNNs , categorized as follows:
Relation-Based HGNNs.
RGCN: Introduces a relational graph convolutional network that assigns weights to each relation type, enabling the learning of structural information in heterogeneous graphs.
RGAT: Applies attention mechanisms to each relation type in relational graphs, effectively integrating information in heterogeneous graphs.
HGT: Utilizes a modified transformer architecture that considers various node and edge types in heterogeneous graphs to learn complex relationships.
HetSANN: Applies attention mechanisms to learn structural information between nodes in heterogeneous graphs without relying on metapaths.
Simple-HGN: Provides efficient and effective node embeddings in heterogeneous graphs through a simplified structure.
Metapath-Based HGNNs.
MP2Vec: Proposes a node embedding technique in heterogeneous networks using metapaths to effectively capture diverse relationships.
HAN: Develops a hierarchical attention mechanism-based model to learn the importance of nodes in heterogeneous graphs using metapaths.
MAGNN: Enhances node embedding performance in heterogeneous graphs by leveraging intermediate node information within metapaths.
HeCo: Combines self-supervised learning and contrastive learning in heterogeneous graphs to improve node representation learning performance.
MHNF: A framework that automatically extracts metapaths in heterogeneous networks and utilizes them for node embedding learning.
HGMAE: Utilizes masked autoencoders to learn robust node representations in heterogeneous graphs.
HMSG: Integrates multi-scale structural information in heterogeneous graphs to enhance node representation learning performance.
MECCH: Improves node embedding performance in heterogeneous graphs by leveraging contextual information of metapaths.
C. Stability of Performance Across Metapaths
Tables 11, 12, and 13 present performance metrics for the top three models and MF2Vec across the DBLP, ACM, and IMDB datasets, respectively, based on a single metapath. Across all datasets, MF2Vec consistently outperforms other models, achieving the highest Mi-F1, Ma-F1, and AUC scores for all metapaths. On the DBLP dataset (Table 11), MF2Vec excels across all metrics. Similarly, on the ACM dataset (Table 12), it achieves the best performance for both PAP and PSP metapaths. On the IMDB dataset (Table 13), MF2Vec again demonstrates its robustness and consistency, achieving the highest scores across all metapaths.
DBLP | ||||||||||||
Metapath | APTPA | APCPA | APA | ALL | ||||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.670 | 0.685 | 0.912 | 0.918 | 0.925 | 0.988 | 0.387 | 0.434 | 0.683 | 0.910 | 0.921 | 0.987 |
HMSG | 0.604 | 0.498 | 0.843 | 0.889 | 0.887 | 0.990 | 0.345 | 0.265 | 0.563 | 0.929 | 0.925 | 0.988 |
MECCH | 0.825 | 0.819 | 0.946 | 0.936 | 0.933 | 0.993 | 0.527 | 0.529 | 0.746 | 0.953 | 0.951 | 0.994 |
MF2Vec | 0.904 | 0.907 | 0.986 | 0.958 | 0.956 | 0.990 | 0.667 | 0.670 | 0.861 | 0.963 | 0.962 | 0.995 |
ACM | |||||||||
Metapath | PAP | PSP | ALL | ||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.6064 | 0.6602 | 0.8754 | 0.7298 | 0.7649 | 0.8575 | 0.8095 | 0.8254 | 0.9480 |
HMSG | 0.6608 | 0.6024 | 0.8598 | 0.7656 | 0.7128 | 0.8302 | 0.8204 | 0.8160 | 0.9151 |
MECCH | 0.6683 | 0.6405 | 0.8431 | 0.7531 | 0.7067 | 0.8440 | 0.7980 | 0.7922 | 0.9194 |
MF2Vec | 0.8080 | 0.8013 | 0.9322 | 0.7631 | 0.7211 | 0.8436 | 0.8653 | 0.8610 | 0.9528 |
IMDB | |||||||||
Metapath | MAM | MDM | ALL | ||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.5241 | 0.5276 | 0.7072 | 0.5057 | 0.5180 | 0.6971 | 0.5685 | 0.5731 | 0.7493 |
HMSG | 0.4717 | 0.4587 | 0.6213 | 0.4226 | 0.3634 | 0.5554 | 0.4767 | 0.4357 | 0.6306 |
MECCH | 0.4149 | 0.3169 | 0.5887 | 0.4029 | 0.3601 | 0.5603 | 0.4639 | 0.4404 | 0.6449 |
MF2Vec | 0.5875 | 0.5733 | 0.7180 | 0.5851 | 0.5565 | 0.7142 | 0.6211 | 0.6117 | 0.7709 |
Model | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
GCN (Ours) | 0.8947 | 0.8200 | 0.8261 | 0.9074 | 0.8283 | 0.8366 | 0.9958 | 0.9631 | 0.9623 | 0.9528 | 0.8653 | 0.8610 | 0.7709 | 0.6211 | 0.6117 | 0.8701 | 0.7429 | 0.6991 |
Edge GAT | 0.8624 | 0.8041 | 0.7941 | 0.8629 | 0.8086 | 0.8180 | 0.9900.002 | 0.9280.007 | 0.9210.007 | 0.9500.005 | 0.8540.008 | 0.8490.010 | 0.7450.013 | 0.5790.023 | 0.5580.025 | 0.7950.010 | 0.6810.010 | 0.6140.026 |
GraphSAGE | 0.8296 | 0.7701 | 0.7771 | 0.8028 | 0.7506 | 0.7719 | 0.9830.005 | 0.9020.020 | 0.8960.022 | 0.9380.006 | 0.8160.009 | 0.8090.009 | 0.7020.018 | 0.5310.029 | 0.5010.031 | 0.7340.018 | 0.6400.021 | 0.4710.016 |
We evaluate the performance of these models across three downstream tasks: node classification, link prediction, and node clustering.
D. Ablation Study
D.1. Effect of Temperature ()
From Figure 10, we observe the sensitivity of the Gumbel-Softmax in calculating the probability across all classes with varying . According to the performance metrics, the model’s performance exhibits minor variations in Movielens, DBLP, and Yelp, and changes within a 2% range in Movielens and Freebase. The results indicate that the sensitivity of the model to different values of is relatively stable across datasets, with minor fluctuations in the Macro F1 and Micro F1 scores.
D.2. Graph Method
Table 14 compares GCN, EdgeGAT, and GraphSAGE to evaluate the impact of facet features as edge weights. GCN outperforms across all datasets, achieving AUC and F1 scores above 0.99 on DBLP and ACM by effectively leveraging edge weights to capture local and global relationships. While EdgeGAT models relationship importance using attention, its performance is less consistent, and GraphSAGE shows lower results due to its limited capacity to represent multi-facet relationships despite its efficiency. These findings highlight the significance of edge weights in facet-based graph learning.
F. Algorithms for Random Walk and Sub-graph Generation
This section describes the algorithms used for random walk and sub-graph generation in heterogeneous networks. The random walk algorithm explores paths within the graph, while the sub-graph generation algorithm utilizes these paths to construct sub-graphs based on specific node types.
F.1. Random Walk Algorithm
The first algorithm 2 performs a random walk on a network. It generates a path of a specified length starting from a given node. The path is built by selecting a random neighbor at each step. If a neighbor cannot be found or if the end node of the path does not match the desired node type, the algorithm returns None.
F.2. Sub-graph Generation Using Random Walk Paths
The second algorithm 3 generates a sub-graph from the network using multiple random walks. For each node, several random walks are performed to collect paths. Unique node pairs are recorded to avoid duplicates. The sub-graph is constructed by adding nodes and edges based on the collected paths.

### v3

@@
 These results highlight its capability to analyze complex networks comprehensively.
-The implementation of MF2Vec is publicly available at https://anonymous.4open.science/r/MF2Vec-6ABC.
+The implementation of MF2Vec is publicly available at https://github.com/kimjongwoo-cell/MF2Vec.
 1. Introduction
@@
 Among these, relation-based HGNNs enhance learning by aggregating messages along edges categorized by relation types, yet they often encounter limitations due to the deep network structures required to process multi-hop relations, leading to over-smoothing and diminished performance (Fu and King, 2024).
+Plot showing Macro-F1 and time complexity comparison between MECCH and MAGNN on DBLP dataset.
 To mitigate these challenges, meta-path-based HGNNs leverage meta-paths to construct homogeneous subgraphs, reducing the need for multi-hop message passing.
@@
 Third, the elimination of manual meta-path definitions and the reduction of computational overhead further bolster scalability.
-Figure 2 demonstrates that even when following the same meta-path, nodes can be grouped based on different facets.
-For example, the paths “New York-L.
-Dicaprio-The Wolf of Wall Street” and “Interstellar-Anne H.-Intern” may simply be grouped as “Movie-Actor-Movie” under meta-path-based methods.
-However, with a multi-facet approach, the path “New York-L.
-Dicaprio-The Wolf of Wall Street” can be considered from the perspective of filming location, while the path “Interstellar-Anne H.-Intern” may reflect similarities in IMDb ratings.
+Schema comparison on IMDB data: (A) predefined node-type schemas vs. (B) dynamic multi-facet path assignment.
+Figure 2 shows that when following the same meta-path, nodes can be grouped based on different facets.
+For example, the paths “New York–L.
+Dicaprio–The Wolf of Wall Street” and “Interstellar–Anne H.–Intern” may simply be grouped as “Movie–Actor–Movie” under meta-path-based methods.
+However, with a multi-facet approach, the path “New York–L.
+Dicaprio–The Wolf of Wall Street” can be considered from the perspective of filming location, while the path “Interstellar–Anne H.–Intern” may reflect similarities in IMDb ratings.
 This approach allows the model to comprehensively capture latent and rich information within the graph, effectively overcoming the limitations of traditional meta-path-based methods.
@@
 These facets characterize the relationships between endpoint nodes and compute path weights that reflect their importance for node connections.
-Finally, a homogeneous network is constructed by integrating same-typed nodes and fed into a Graph Convolutional Network (GCN) (Kipf and Welling, 2016), which aggregates multi-facet vectors along the paths to generate latent node representations.
+Finally, a homogeneous network is constructed by integrating same-typed nodes and fed into a GNN, which aggregates multi-facet vectors along the paths to generate latent node representations.
 Key contributions of this work are summarized as follows:
 Dynamic Facet Embedding: A novel framework for constructing flexible paths independent of node types.
-Efficiency and Performance Enhancement: Reduction in computational costs and enhanced performance for node classification, node clustering, and link prediction tasks
 Stable Performance: Significant improvement in stability across datasets and tasks by mitigating dependence on fixed meta-paths.
+Efficiency and Performance Enhancement: Reduced computational costs and improved performance for node classification, node clustering, and link prediction tasks.
+2. Preliminary
+Definition 1 (Heterogeneous Graph). A heterogeneous graph is defined by a node type mapping function and an edge type mapping function , where and are sets of node and edge types, respectively, with .
+Definition 2 (Metapath). A metapath is a sequence (or ), representing a composite relation between node types and .
+Definition 3 (Metapath Schema).
+A metapath schema is a sequence of node types connected by edge types .
+This represents the structural pattern of a metapath and illustrates how different node types are related through specific types of relationships defined by the nodes and edges.
+Definition 4 (Multi-facet Node Representation Learning).
+For a graph with nodes , the goal is to learn a multi-facet node embedding for each node .
+Here, is the number of facets, and each facet embedding aims to: 1) preserve network structure, 2) capture various facets of , and 3) represent facet interactions.
+Definition 5 (Multi-facet Path).
+Unlike predefined metapaths, a multi-facet path is extracted without manual specification.
+Intermediate nodes are projected onto multi-facet vectors , forming a multi-facet path .
+This path integrates information from intermediate nodes through multiple facets, updating node representations based on various latent information beyond node types.
+Overall architecture diagram of the MF2Vec model pipeline.
+3. Methodology
+In this section, we describe our approach, as shown in Figure 3.
+We begin by extracting paths between homogeneous end nodes using random walks, similar to traditional metapath methods (Dong et al., 2017; Fu et al., 2017).
+We then project the features of intermediate nodes into multiple facets and aggregate these facets to create multi-facet embedding representations.
+These representations are used as edge features between path-guided neighbors and the target node.
+The model pseudocode is provided in Algorithm 1, and the notation is summarized in Table 1.
+Notation | Explanation |
+-dimensional Euclidean space | |
+A heterogeneous graph with a node set and an edge set | |
+A subgraph with a certain node type set and a certain node type edge set | |
+A node | |
+The embedding of node | |
+-th multi-facet embedding of node | |
+Multi-facet embedding of node | |
+Final embedding of node after -th layer | |
+The set of whole nodes in the path between node and | |
+Intermediate nodes in the path between node and | |
+The embedding of the path between node and | |
+-th multi-facet embedding of the path between node and | |
+Multi-facet embedding of the path between node and | |
+Weight of -th facet in the path |
+3.1. Extracting Paths Using Random Walk
+Given a node type , our method constructs a homogeneous subgraph that identifies direct connections between nodes of type , the same as metapath-based models.
+Instead of relying on metapath schemas such as APA or APCPA, the method generates a path starting from a node of type and applies a random walk, traversing randomly selected adjacent nodes until encountering a type A node.
+This approach captures relationships within a specified path length limit, effectively mapping connections within the nodes of a type without being constrained by predefined metapath patterns.
+Details of the Random Walk algorithm are provided in Algorithm 2.
+3.2. Transforming Single Embeddings into Multi-Facet
+Our model captures inherent information from all node types before generating a facet embedding.
+Node information is exchanged through a GNN warm-up phase, which facilitates the aggregation of neighborhood information and enriches the initial node embeddings with structural and semantic context.
+The node embedding, , is defined during this warm-up phase and serves as the foundation for training the main model.
+The resulting node embedding is then refined into a multi-dimensional embedding to capture diverse information, as detailed in Equation 1.
+Here, denotes the embedding of node for the -th facet, where , is a learnable weight matrix of size for each facet, and is the total number of facets.
+The embeddings for all facets are then concatenated to form a unified multi-facet embedding, , providing a comprehensive representation of the node.
+3.3. Multi-facet Feature Embedding Aggregation
+To construct a multi-facet feature embedding for paths consisting of multiple nodes, we employ an aggregation function that synthesizes information from the nodes along the path.
+The mathematical formulation for this aggregation is given by Equation 2.
+In the given equation, represents the aggregated multi-facet feature embedding for a connection between nodes and .
+The set consists of intermediate nodes between and , and is an intermediate node within the path.
+The function signifies an aggregation operation using mean pooling.
+3.4. Discriminative Facet Consideration using Gumbel Softmax
+We utilize the Gumbel Softmax (Jang et al., 2016) function to introduce randomness and uncertainty into facet selection, which helps the model explore diverse facet influences in complex relationships.
+This approach enables us to capture nuanced facet interactions and dependencies, making our model more adaptable and robust for multi-facet data.
+Even if the random walk algorithm extracts the paths that are little related to the end nodes, we mitigate their impact by assigning higher weights to the most crucial components using Gumbel Softmax.
+In the given equation, represents Gumbel Softmax. is the final embedding for that combines these probabilities, and is a matrix with dimensions .
+3.5. Subgraph Construction
+Path-based HGNNs transform the input graph into a new homogeneous subgraph for each node and path, where varies by model.
+For example, in HAN, represents the metapath--guided neighborhood, while MAGNN defines it as instances of metapath- incident to the node.
+MECCH constructs subgraphs using most metapaths as illustrated in Figure 4.
+Our approach, however, avoids the need for multiple metapaths by leveraging multi-facet information from a single path between nodes.
+This results in a subgraph that provides richer information than traditional methods.
+The MF2Vec model efficiently learns complex node relationships by constructing subgraphs incorporating multi-dimensional edge features which are multi-facet embedding for a more comprehensive node representation.
+Details are provided in Algorithm 3.
+Comparison of HGNN aggregation strategies: HAN, MAGNN, MHNF, MECCH, and MF2Vec for a green target node.
+3.6. Message Passing
+To extract node embeddings, we use GCN, considering node interactions and incorporating path facets as edge features. The graph node embedding is computed as follows:
+In this equation, represents the activation function, which is ELU, and BN represents the batch normalization.
+The denotes the path facet feature between node and its neighbor , which is utilized as an edge feature.
+3.7. Training Losses
+We jointly train our model for node classification and link prediction.
+For node classification, we adopt the cross-entropy loss , where is the set of labeled nodes, is the true label, and is the predicted probability.
+For link prediction, we use a contrastive loss , where and are the positive and negative node pairs, respectively, and denotes the sigmoid function.
+These two objectives enable the model to learn both node-level classification and structural link prediction simultaneously.
+4. Experiment
+In this part, we assess the performance and practicality of our introduced method through tests on link prediction, node classification, and node clustering tasks across five diverse graph datasets.
+Datasets.
+We evaluate our model on six heterogeneous graph datasets: DBLP, ACM, IMDB, and Freebase for node classification and clustering, and Yelp and MovieLens for link prediction.
+Dataset details are summarized in Table 2.
+Dataset | # Nodes (# Types) | # Edges | Target (# Labels) | Task |
+Yelp | 31,081 (3) | 411,263 | User-Business | LP |
+Movielens | 2,672 (4) | 234,695 | User-Movie | LP |
+DBLP | 26,128 (4) | 119,783 | Author (4) | NC, CL |
+ACM | 11,246 (3) | 17,426 | Paper (3) | NC, CL |
+IMDB | 11,519 (3) | 17,009 | Movie (3) | NC, CL |
+Freebase | 43,854 (4) | 75,517 | Movie (3) | NC, CL |
+Baselines.
+We compare MF2Vec against SOTA heterogeneous graph neural networks (HGNNs), categorized as follows: (a) Relation-Based HGNNs, including RGCN, RGAT, HGT, HetSANN, and Simple-HGN; and (b) Metapath-Based HGNNs, comprising MP2Vec, HAN, MAGNN, HeCo, MHNF, HGMAE HMSG, and MECCH.
+Graph Construction Settings.
+Following Algorithm 2, we ensured that the node types matched the target node and that the path length was at most 5.
+For each target node, we performed up to 1000 trials.
+The number of negative samples was set to 1.
+Experimental Settings.
+To ensure a fair comparison among all HGNNs, the node embedding dimension was fixed to 64.
+The learning rate was set to , weight decay to , and the dropout rate to 0.5.
+For each baseline model, hyperparameters such as batch size, attention heads, dropout rate, learning rate, and epochs—excluding the node embedding dimension—followed the default settings specified in the original papers.
+Training was conducted using the Adam optimizer with early stopping (patience of 20 epochs) to prevent overfitting.
+Specifically, for MF2Vec, the temperature () was set to 0.5, and the number of facets () was set to 5.
+To demonstrate generalization, the experiments were conducted using five seed values (1, 10, 100, 1000, and 10000), with the data split into train (80%), validation (10%), and test (10%).
+The model seeds were fixed during the experiments, and the average and standard deviation of the results were calculated for all seeds.
+Dataset | DBLP | ACM | IMDB | Freebase | |||||||||
+Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
+Relation-based | RGCN | 0.8580.004 | 0.8420.005 | 0.8370.005 | 0.8100.012 | 0.7690.007 | 0.7620.011 | 0.5330.024 | 0.4310.027 | 0.3370.043 | 0.5740.049 | 0.5080.057 | 0.4080.078 |
+RGAT | 0.8490.003 | 0.8290.004 | 0.8230.004 | 0.8030.021 | 0.7610.023 | 0.7490.025 | 0.5580.011 | 0.4240.015 | 0.4090.016 | 0.6220.027 | 0.5530.033 | 0.4870.037 | |
+HGT | 0.9260.017 | 0.8890.026 | 0.8990.022 | 0.7790.016 | 0.7410.016 | 0.7090.024 | 0.5400.019 | 0.3960.024 | 0.3860.022 | 0.5780.030 | 0.4820.034 | 0.4270.037 | |
+HetSANN | 0.9500.006 | 0.9310.007 | 0.9250.008 | 0.7800.023 | 0.7360.025 | 0.7110.027 | 0.6130.006 | 0.4870.010 | 0.4840.007 | 0.6460.021 | 0.5750.027 | 0.5210.030 | |
+Simple-HGN | 0.9470.011 | 0.9280.011 | 0.9200.015 | 0.8400.037 | 0.8080.033 | 0.7930.045 | 0.6120.035 | 0.4970.038 | 0.4810.045 | 0.6710.014 | 0.6300.017 | 0.5480.031 | |
+Metapath-based | MP2Vec | 0.5860.020 | 0.3990.025 | 0.3720.026 | 0.7310.010 | 0.6630.010 | 0.6330.014 | 0.5110.011 | 0.3970.016 | 0.3780.013 | 0.6050.019 | 0.6030.022 | 0.5060.030 |
+HAN | 0.9370.003 | 0.9190.003 | 0.9090.004 | 0.7850.011 | 0.7620.009 | 0.7290.014 | 0.5500.020 | 0.4260.021 | 0.3840.034 | 0.6650.011 | 0.6430.014 | 0.5170.033 | |
+MAGNN | 0.9720.003 | 0.9260.003 | 0.9150.003 | 0.9450.014 | 0.8060.031 | 0.7770.049 | 0.7390.013 | 0.5340.032 | 0.4920.070 | 0.7440.015 | 0.6740.019 | 0.5180.063 | |
+HeCo | 0.9780.003 | 0.8940.013 | 0.8860.014 | 0.8310.019 | 0.7450.012 | 0.7030.013 | 0.5440.006 | 0.3850.014 | 0.3600.004 | 0.7530.021 | 0.6520.021 | 0.5150.033 | |
+MHNF | 0.9250.004 | 0.8960.006 | 0.8860.006 | 0.8830.009 | 0.7600.012 | 0.7110.016 | 0.5300.042 | 0.4080.042 | 0.3240.071 | 0.5120.013 | 0.4510.014 | 0.2650.053 | |
+HGMAE | 0.9480.018 | 0.8740.121 | 0.8290.084 | 0.8120.011 | 0.7320.012 | 0.7360.019 | 0.6140.019 | 0.4270.016 | 0.3530.033 | 0.7270.030 | 0.6220.030 | 0.4780.042 | |
+HMSG | 0.9780.003 | 0.9010.020 | 0.8940.021 | 0.9230.013 | 0.7950.019 | 0.7850.020 | 0.6470.021 | 0.4720.017 | 0.4610.022 | 0.7440.026 | 0.6790.040 | 0.5310.038 | |
+MECCH | 0.9910.133 | 0.9410.165 | 0.9370.176 | 0.9160.133 | 0.7900.165 | 0.7930.176 | 0.6570.133 | 0.5120.165 | 0.5020.176 | 0.7730.133 | 0.6480.165 | 0.5880.176 | |
+MF2Vec | 0.9920.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
+4.1. Downstream Task
+4.1.1. Node Classification.
+In our node classification experiments, we convert nodes into low-dimensional vector embeddings and apply a Softmax function to obtain probability distributions over label categories.
+We evaluate the model using Micro-F1, Macro-F1, and AUC metrics, providing a balanced assessment of classification performance.
+These metrics are used to compare our model, MF2Vec, against baseline models on the DBLP, ACM, IMDB and Freebase datasets, with results presented in Table LABEL:table_1.
+Our approach consistently outperforms baseline models across all metrics, demonstrating the effectiveness of MF2Vec.
+The significant performance improvement, particularly on the IMDB dataset, highlights the model’s ability to capture complex relational data through multi-facet vectors.
+This suggests that MF2Vec effectively reveals nuanced relationships within heterogeneous information networks (HINs), enhancing understanding beyond node types.
+4.1.2. Link Prediction.
+In Movielens and Yelp, link prediction is framed as a binary classification task to determine the presence of an edge in the original graph.
+During training, validation, and testing, negative edges are created by replacing the destination node in positive edges with a randomly chosen node that is not connected to the source node.
+The link probability is computed using the dot product of node embeddings: , where is the sigmoid function.
+Table 4 presents the AUC, Micro-F1, and Macro-F1 scores for our model and baselines on Yelp and Movielens.
+MF2Vec performs competitively on Yelp (against MECCH) and outperforms all baselines on MovieLens.
+These results highlight the advantage of using multi-facet vectors derived from node paths for link prediction, and demonstrate the efficacy of our model in updating embeddings without relying on node types.
+Dataset | Yelp | Movielens | ||||
+Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
+RGCN | 0.6350.141 | 0.6210.103 | 0.6140.134 | 0.7410.003 | 0.6640.003 | 0.6640.003 |
+RGAT | 0.6970.104 | 0.6650.081 | 0.6550.086 | 0.6870.115 | 0.6360.101 | 0.6300.104 |
+HGT | 0.8320.004 | 0.7710.004 | 0.7700.005 | 0.7550.136 | 0.7110.116 | 0.7110.115 |
+HetSANN | 0.7030.017 | 0.6890.013 | 0.6880.014 | 0.7740.013 | 0.7720.013 | 0.7710.013 |
+Simple-HGN | 0.8240.016 | 0.7570.018 | 0.7570.018 | 0.8200.007 | 0.7460.012 | 0.7450.015 |
+MP2Vec | 0.7180.003 | 0.6580.002 | 0.6640.003 | 0.7980.004 | 0.6620.004 | 0.6690.002 |
+HAN | 0.6270.102 | 0.5280.076 | 0.3990.140 | 0.8220.004 | 0.6420.050 | 0.6040.072 |
+MAGNN | 0.8350.005 | 0.7730.002 | 0.7740.003 | 0.8580.010 | 0.7760.008 | 0.7810.009 |
+GraphHINGE | 0.8390.001 | 0.7730.001 | 0.7660.002 | 0.8580.004 | 0.7750.003 | 0.7830.004 |
+HeCo | 0.8380.003 | 0.7670.004 | 0.7550.002 | 0.8390.003 | 0.7550.004 | 0.7710.005 |
+HMSG | 0.8510.003 | 0.7840.002 | 0.8030.004 | 0.8140.006 | 0.7440.006 | 0.7530.013 |
+MECCH | 0.8670.012 | 0.7820.013 | 0.7740.015 | 0.7520.006 | 0.6900.015 | 0.6340.040 |
+MF2Vec | 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 |
+4.1.3. Node Clustering.
+We carry out node clustering experiments using the DBLP, ACM, IMDB, and Freebase datasets to assess the embedding quality produced by MF2Vec.
+In these experiments, similar to node classification, we utilize the embedding vectors of target nodes from the testing set as input for the K-Means model and employ NMI and ARI metrics to gauge performance.
+As presented in Table 5, our model surpasses the baseline models, except for MHNF (Sun et al., 2022) in ARI on the IMDB dataset, where MF2Vec closely follows it by a small margin, demonstrating our model’s superior ability to generate effective node representations based on multi-facet vectors for node clustering.
+Dataset | DBLP | ACM | IMDB | Freebase | ||||
+Model | NMI | ARI | NMI | ARI | NMI | ARI | NMI | ARI |
+RGCN | 0.0168 | 0.0670 | 0.3612 | 0.2809 | 0.0088 | 0.0056 | 0.0228 | 0.0443 |
+RGAT | 0.4435 | 0.4097 | 0.3976 | 0.3140 | 0.0477 | 0.0143 | 0.0327 | 0.0387 |
+HGT | 0.7722 | 0.8250 | 0.5064 | 0.5335 | 0.0433 | 0.0459 | 0.0801 | 0.1075 |
+HetSANN | 0.752 | 0.7723 | 0.5274 | 0.5263 | 0.0865 | 0.043 | 0.1054 | 0.0967 |
+Simple-HGN | 0.7707 | 0.6405 | 0.5309 | 0.5237 | 0.1410 | 0.1180 | 0.1541 | 0.1804 |
+MP2Vec | 0.0519 | 0.0330 | 0.1034 | 0.0641 | 0.0065 | 0.0096 | 0.0145 | 0.0286 |
+HAN | 0.3533 | 0.1845 | 0.3036 | 0.3528 | 0.0057 | 0.0015 | 0.0211 | 0.0217 |
+MAGNN | 0.7964 | 0.8193 | 0.5125 | 0.5317 | 0.1075 | 0.1199 | 0.1872 | 0.1477 |
+HeCo | 0.7477 | 0.7872 | 0.5394 | 0.5454 | 0.0049 | 0.0028 | 0.0805 | 0.0793 |
+MHNF | 0.7197 | 0.7498 | 0.5385 | 0.5274 | 0.1443 | 0.1467 | 0.1881 | 0.1697 |
+HGMAE | 0.7076 | 0.7375 | 0.5315 | 0.5321 | 0.1264 | 0.1345 | 0.2401 | 0.2485 |
+HMSG | 0.7278 | 0.7767 | 0.4475 | 0.3630 | 0.0335 | 0.0326 | 0.1257 | 0.1213 |
+MECCH | 0.8312 | 0.8739 | 0.4202 | 0.3513 | 0.0070 | 0.0056 | 0.2612 | 0.3127 |
+MF2Vec | 0.8606 | 0.8951 | 0.5714 | 0.5790 | 0.1459 | 0.1448 | 0.3083 | 0.3729 |
+4.2. Stability under Meta-path Variations
+This section empirically evaluates whether MF2Vec can achieve consistent and robust performance across various settings by effectively leveraging multi-facet information shared among all nodes, regardless of the presence or absence of node type-based metapaths.
+All experiments are conducted on three datasets, and the relative advantages of MF2Vec are assessed through comparison with representative metapath-based GNNs (MAGNN, HMSG, MECCH).
+The details and interpretations of each experiment are as follows.
+4.2.1. Consistency Across Metapath Types
+Table 6 presents the performance variability of each model when trained and evaluated on different metapath types (e.g., APA, APCPA, etc.).
+For each dataset, we measure the standard deviation (Std) of AUC, Mi-F1, and Ma-F1 scores, and find that MF2Vec consistently achieves the lowest standard deviation, indicating the highest stability.
+This result suggests that MF2Vec does not rely on specific metapaths defined by node types, but instead learns consistent and reliable representations by leveraging shared facet information among nodes.
+Dataset | Metric | MAGNN | HMSG | MECCH | MF2Vec |
+DBLP | AUC | 0.130 | 0.177 | 0.107 | 0.060 |
+Mi-F1 | 0.217 | 0.222 | 0.173 | 0.127 | |
+Ma-F1 | 0.201 | 0.257 | 0.170 | 0.125 | |
+ACM | AUC | 0.038 | 0.039 | 0.034 | 0.032 |
+Mi-F1 | 0.078 | 0.058 | 0.051 | 0.033 | |
+Ma-F1 | 0.052 | 0.075 | 0.062 | 0.049 | |
+IMDB | AUC | 0.034 | 0.039 | 0.041 | 0.025 |
+Mi-F1 | 0.020 | 0.023 | 0.049 | 0.008 | |
+Ma-F1 | 0.015 | 0.046 | 0.079 | 0.010 |
+4.2.2. Generalization Across Unseen Metapaths
+Figure 5 compares the generalization performance of each model when trained on a single metapath and tested on different, unseen metapaths.
+MF2Vec consistently achieves the highest generalization performance even under these unseen conditions, and the performance drop compared to using all metapaths together is significantly smaller than that of other models.
+This indicates that MF2Vec is less dependent on specific metapath schemas and is able to extract more transferable and robust representations.
+These results confirm that the ability to select facets independently of explicit path structures is crucial for achieving stable and generalized performance.
+Comparison of cross-metapath generalization performance across models. MF2Vec shows smallest performance drop.
+Dataset | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
+Method | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
+No.
+| 0.8130.003 | 0.7430.002 | 0.7520.004 | 0.8320.003 | 0.7540.005 | 0.7430.013 | 0.9740.004 | 0.8680.012 | 0.8540.012 | 0.9120.014 | 0.7880.015 | 0.7590.025 | 0.531 0.022 | 0.378 0.005 | 0.232 0.036 | 0.7310.011 | 0.6460.021 | 0.4750.015 |
+Random.
+| 0.8650.008 | 0.7920.004 | 0.8030.009 | 0.8690.006 | 0.7930.007 | 0.8020.004 | 0.9850.004 | 0.9320.007 | 0.9300.005 | 0.9450.005 | 0.8230.008 | 0.8120.006 | 0.7480.008 | 0.5840.003 | 0.5690.007 | 0.8240.008 | 0.6910.008 | 0.6260.007 |
+STE.
+| 0.8670.018 | 0.7970.003 | 0.8080.015 | 0.8710.006 | 0.7950.006 | 0.8030.005 | 0.9890.003 | 0.9350.009 | 0.9310.01 | 0.9490.009 | 0.8240.013 | 0.8130.012 | 0.7510.005 | 0.5880.004 | 0.5710.003 | 0.8250.020 | 0.6960.017 | 0.6280.042 |
+Gumble.
+| 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
+4.3. Time Complexity
+Model | Time Complexity |
+HetSANN | |
+SimpleHGN | |
+MAGNN | |
+HAN | |
+MECCH | |
+MF2Vec |
+The proposed model’s time complexity is , where is the number of nodes, is the average neighbors, is the number of facets, and is the embedding dimension.
+Attention adds , and graph convolution contributes , resulting in a total complexity of .
+In Table 8, models like MECCH (Fu and King, 2024), HAN (Wang et al., 2019), and MAGNN (Fu et al., 2020) rely on metapaths, incurring higher costs due to growth with metapath length ().
+MF2Vec avoids this by leveraging facets (), leading to more efficient computation, satisfying .Additionally, we compare MF2Vec with relation-based methods like HetSANN (Hong et al., 2020) and SimpleHGN (Lv et al., 2021), which also avoid metapaths by focusing on edge types.
+Figure 6 compares training times and Macro F1 scores for MF2Vec and other models on DBLP, IMDB, and Movielens.
+MAGNN (Fu et al., 2020) and MP2Vec (Dong et al., 2017) are excluded from Movielens due to their excessive training times.
+MF2Vec achieves the best performance on node classification and link prediction tasks, with significantly lower training times.
+This efficiency arises from its dynamic use of path characteristics, avoiding the high costs of aggregating multiple paths as seen in MAGNN and MP2Vec, ensuring better scalability and faster convergence without compromising performance.
+Training time (log seconds) vs Macro F1 across three datasets. MF2Vec achieves best tradeoff with faster convergence.
+4.4. Effect of Facet Selection Strategy
+Since MF2Vec selects the optimal facet for each path in a learnable manner, we conducted an additional experiment to validate the effectiveness of its facet selection strategy.
+Table 7 compares various methods for selecting facet information as edge weights.
+Our GCN-based results show that the proposed method, which leverages Gumbel-Softmax and attention mechanisms, consistently outperforms all alternatives across datasets.
+When using random facet selection, performance decreases, and omitting facet information entirely leads to the lowest results.
+We also evaluated the STE (Straight-Through Estimator)-based approach (Bengio et al., 2013), which enables probabilistic gradient propagation during facet selection.
+However, this method performs slightly worse than our proposed approach.
+These findings highlight that facet-related information is crucial for effective graph learning, and our method excels at capturing complex structures and hidden relationships.
+4.5. Visualization
+To evaluate embedding quality, we visualize node representations from various HGNNs using t-SNE (Van der Maaten and Hinton, 2008).
+Figure 7 shows t-SNE plots for HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MECCH (Fu and King, 2024), and MF2Vec, with colors representing author classes.
+MF2Vec demonstrates superior clustering within classes and clear decision boundaries, outperforming the more scattered intra-class distributions and weaker inter-class separations of other HGNNs.
+Scatter plot showing the learned embeddings of nodes in DBLP, colored by label, demonstrating clear clustering by class.
+4.6. Sensitivity to Random Walk
+4.6.1. Effect of Random Walk Length.
+In Table 9, increasing the random walk length potentially introduces more noise into the sampled paths, but the performance of MF2Vec remains stable across all tested values.
+Changes in AUC are minor, demonstrating that MF2Vec is insensitive to longer or noisier random walks.
+Based on these results, a length of 5 was used for experiments.
+Dataset | 3 | 5 | 7 | 9 | |
+DBLP | 0.990 0.008 | 0.992 0.002 | 0.988 0.010 | 0.982 0.009 | 0.010 |
+ACM | 0.939 0.010 | 0.954 0.008 | 0.953 0.007 | 0.951 0.011 | 0.015 |
+IMDB | 0.748 0.025 | 0.768 0.033 | 0.772 0.021 | 0.779 0.014 | 0.031 |
+4.6.2. Effect of Random Walk Quality.
+We simulate sparse environments by randomly removing 0%, 10% and 20% of node edges, then evaluate AUC and its standard deviation.
+As shown in Figure 8, MF2Vec shows the least performance variation, maintaining robust results despite increased noise and missing data.
+This demonstrates that MF2Vec leverages shared facet information to generalize well in incomplete or noisy settings.
+Line chart comparing robustness of MF2Vec and other models when 0, 10, or 20 percent of edges are randomly removed. MF2Vec shows consistently higher performance across settings.
+4.7. Analysis of Facet Balance
+To further analyze the balance of facet distributions across node types, we visualized a facet-level attention heatmap as shown in Figure 9.
+Each column represents a specific node and each row corresponds to a facet.
+Darker colors indicate facets that play a more significant role in the node’s connections.
+For example, the paper node “P1612” mainly focuses on Facet 1, whereas the author node “A3958” distributes attention relatively evenly across all facets.
+This analysis confirms that some nodes concentrate on specific facets while others distribute attention more uniformly, with no strong bias overall.
+This visually validates the importance of modeling shared facets in heterogeneous graphs.
+A heatmap showing facet-level attention across nodes in the DBLP dataset.
+Columns correspond to sampled nodes of different types, rows correspond to facets, and darker cells represent more influential facets.
+4.8. Case Study
+A path in the IMDB graph with nodes annotated by attributes such as rating, country, and language. All nodes share the language English, showing semantic grouping by facet.
+Most graph datasets represent nodes by unique IDs, making it difficult to interpret learned embeddings in terms of real-world attributes.
+To address this, we conducted a qualitative case study to investigate whether the learned facets actually correspond to meaningful attributes.
+Based on the IMDB text data provided in MAGNN, we used SBERT to generate node embeddings and separated them into five explicit attributes shared by IMDB node types: rating, genre, country, language, and keywords.
+In this setting, we analyzed which attributes were shared among the intermediate nodes along a given path from a start node to an end node.
+For example, along the path “Horrible Bosses 2 → Penelope Wilton → All the Pretty Horses → Kevin Peter Hall → The Sorcerer’s Apprentice,” all nodes shared the attributes country (USA, UK), rating (around 6), and language (English).
+Notably, Gumbel-Softmax identified “language (English)” as the dominant facet.
+This shows that the connections in this path are not just simple Movie–Actor–Movie–Actor–Movie links, but are semantically grouped by language, thus demonstrating that MF2Vec learns interpretable and meaningful facets rather than mere latent ID embeddings.
+4.9. Ablation Study
+4.9.1. Effect of Temperature ()
+From Figure 11, we observe the sensitivity of the Gumbel-Softmax in calculating the probability across all classes with varying .
+According to the performance metrics, the model’s performance exhibits minor variations in Movielens, DBLP, and Yelp, and changes within a 2% range in Movielens and Freebase.
+The results indicate that the sensitivity of the model to different values of is relatively stable across datasets, with minor fluctuations in the Macro F1 and Micro F1 scores.
+A line plot showing MF2Vec performance across six datasets at different temperature values in Gumbel-Softmax. Performance varies slightly but remains stable overall.
+4.9.2. Effect of Train Ratio Settings.
+As shown in Figure 12, MF2Vec consistently achieves the highest performance across varying training set ratios, demonstrating its robustness and adaptability.
+While the performance of all models declines as the training ratio decreases, MF2Vec remains superior in most cases and retains competitiveness in the rest.
+At the lowest ratio of 0.2, it occasionally shows comparable performance to other models.
+Only models achieving scores above a certain threshold on each dataset are displayed to ensure clearer visualization.
+These results highlight MF2Vec’s effectiveness in handling sparse training data.
+A line chart showing classification performance of MF2Vec compared to baselines when varying the training set ratio from 0.8 to 0.2. MF2Vec consistently outperforms baselines.
+4.9.3. Effect of Number of Facets.
+Figure 13 shows MF2Vec’s performance across six datasets with varying facet counts.
+The performance difference relative to “” highlights the advantage of multiple facets.
+Except for IMDB, optimal results are achieved with five facets, consistently outperforming the single-facet case by effectively leveraging node and path information.
+Beyond five facets, performance declines, likely due to reduced feature grouping efficiency.
+A plot showing MF2Vec’s performance trend as the number of facets increases. Best performance occurs at five facets, with degradation beyond that point.
 5. Related Work
-2.1. Heterogeneous Graph Neural Network (HGNN)
+5.1. Heterogeneous Graph Neural Network
 Graph neural networks (GNNs) effectively handle graph-structured data by learning low-dimensional node representations (Defferrard et al., 2016; Kipf and Welling, 2016; Hamilton et al., 2017; Velickovic et al., 2017).
@@
 We extend these approaches to heterogeneous graphs by transforming meta-path schemas into diverse facets, linking terminal node facets for richer interaction understanding.
-3. Preliminary
-This section provides definitions of terminologies used in the current study. The notations and their explanations are summarized in Table 1.
-Definition 1 (Heterogeneous Graph).
-A heterogeneous graph is characterized by a node type mapping function and an edge type mapping function , where and are sets of node and edge types, respectively, with .
-Definition 2 (Metapath). A metapath is a sequence (or ), representing a composite relation between node types and .
-Definition 3 (Metapath Schema).
-A metapath schema is a sequence of node types connected by edge types .
-This schema represents the structural pattern of a metapath and illustrates how different node types are related through specific types of relationships defined by the nodes and edges.
-Definition 4 (Multi-facet Node Representation Learning).
-For a graph with nodes , the goal is to learn a multi-facet node embedding for each node .
-Here, is the number of facets, and each facet embedding aims to: 1) preserve network structure, 2) capture various facets of , and 3) represent facet interactions.
-Definition 5 (Multi-facet Path).
-Unlike predefined metapaths, a multi-facet path is extracted without manual specification.
-Intermediate nodes are projected onto multi-facet vectors , forming a multi-facet path .
-This path integrates information from intermediate nodes through multiple facets, updating node representations based on various latent information beyond node types.
-4. Methodology
-In this section, we describe our approach, as shown in Figure 3.
-We begin by extracting paths between homogeneous end nodes using random walks, similar to traditional metapath methods (Dong et al., 2017; Fu et al., 2017).
-We then project the features of intermediate nodes into multiple facets and aggregate these facets to create multi-facet embedding representations.
-These representations are used as edge features between path-guided neighbors and the target node.
-The model pseudocode is provided in Algorithm 1, and the notations are summarized in Table 1.
-Notation | Explanation |
--dimensional Euclidean space | |
-A heterogeneous graph with a node set and an edge set | |
-A subgraph with a certain node type set and a certain node type edge set | |
-A node | |
-The embedding of node | |
--th multi-facet embedding of node | |
-Multi-facet embedding of node | |
-Final embedding of node after -th layer | |
-The set of whole nodes in the path between node and | |
-Intermediate nodes in the path between node and | |
-The embedding of the path between node and | |
--th multi-facet embedding of the path between node and | |
-Multi-facet embedding of the path between node and | |
-Weight of -th facet in the path |
-4.1. Extracting Paths Using Random Walk
-Given a node type , our method constructs a homogeneous subgraph that identifies direct connections between nodes of type , the same as traditional metapath-based models.
-Instead of relying on metapath schemas such as APA or APCPA, the method generates a path starting from a node of type and applies a random walk, traversing randomly selected adjacent nodes until encountering a type A node.
-This approach captures relationships within a specified path length limit, effectively mapping connections within the nodes of a type without being constrained by predefined metapath patterns.
-Details of the Random Walk algorithm are provided in Algorithm 2.
-4.2. Transforming Single Embeddings into Multi-Facet
-Our model captures inherent information from all node types before generating a facet embedding.
-Initially, node information is exchanged through a GNN warm-up phase, which facilitates the aggregation of neighborhood information and enriches the initial node embeddings with structural and semantic context.
-The node embedding, , is defined during this warm-up phase and serves as the foundation for training the main model.
-The resulting node embedding is then refined into a multi-dimensional embedding to capture diverse information, as detailed in Equation 1.
-Here, represents the multi-facet embedding of node within the -th facet, where is the corresponding weight matrix with dimensions for the -th facet, and represents the total number of facets.
-The individual facet embeddings are subsequently concatenated into a unified multi-facet embedding, , providing a comprehensive representation.
-4.3. Multi-facet Feature embedding Aggregation
-To construct a multi-facet feature embedding for paths consisting of multiple nodes, we employ an aggregation function that synthesizes information from the nodes along the path.
-The mathematical formulation for this aggregation is given by Equation 2.
-In the given equation, represents the aggregated multi-facet feature embedding for a connection between nodes and .
-The set consists of intermediate nodes between and , and is an intermediate node within the path.
-The function signifies an aggregation operation using mean pooling.
-4.4. Discriminative Facet Consideration using Gumbel Softmax
-We utilize the Gumbel Softmax (Jang et al., 2016) function to introduce randomness and uncertainty into facet selection, which helps the model explore diverse facet influences in complex relationships.
-The temperature parameter allows us to control the balance between exploration and exploitation.
-This approach enables us to capture nuanced facet interactions and dependencies, making our model more adaptable and robust for multi-facet data.
-Even if the random walk algorithm extracts the paths that are little related to the end nodes, we mitigate their impact by assigning higher weights to the most crucial components using Gumbel softmax.
-As a result, we effectively operate within constrained environments, even without extracting meaningful paths.
-In the given equation, represents Gumbel Softmax. is the final embedding for that combines these probabilities.and is a matrix with dimensions .
-4.5. Subgraph Construction
-Path-based HGNNs transform the input graph into a new homogeneous subgraph for each node and path, where varies by model.
-For example, in HAN, represents the metapath--guided neighborhood, while MAGNN defines it as instances of metapath- incident to the node.
-MECCH constructs subgraphs using most metapaths as illustrated in Figure 4.
-Our approach, however, avoids the need for multiple metapaths by leveraging multi-facet information from a single path between nodes.
-This results in a subgraph that provides richer information than traditional methods.
-The MF2Vec model efficiently learns complex node relationships by constructing subgraphs as described in Algorithm 3 incorporating multi-dimensional edge features which are multi-facet embedding for a more comprehensive node representation.
-4.6. Message Passing
-To extract node embeddings, we utilize GCN.
-In this process, we consider interactions between nodes and utilize path facets as edge features.
-The formula for calculating graph node embeddings is as follows:
-In this equation, represents the activation function, which is ELU, and BN represents the batch normalization.
-The denotes the path facet feature between node and its neighbor , which is utilized as an edge feature.
-4.7. Training Losses
-We train our model for two tasks: node classification and link prediction.
-For node classification, the cross-entropy loss is computed over labeled nodes , measuring the disparity between true labels and predictions :
-For link prediction, where node labels are unavailable, we minimize the contrastive loss using positive and negative node pairs:
-Here, is the sigmoid function, ensuring effective learning for link prediction.
-5. Experiment
-In this part, we assess the performance and practicality of our introduced method through tests on link prediction, node classification, and node clustering tasks across five diverse graph datasets.
-The primary goals of these experiments are to address the following research questions.
-RQ1. How well does MF2Vec perform on downstream tasks?
-RQ2. How stable is the performance of MF2Vec in downstream tasks?
-RQ3. How time-efficient is MF2Vec?
-5.1. Experimental Settings
-5.1.1. Datasets
-We evaluate our model on six heterogeneous graph datasets: DBLP, ACM, IMDB, and Freebase for node classification and clustering, and Yelp and MovieLens for link prediction.
-Experiments include 1000 trials per node with a walk length of 5, covering typical metapaths (e.g., APA, APCPA) using consistent data splits.
-Dataset details are summarized in Table 2.
-Dataset | # Nodes (# Types) | # Edges | Target (# Labels) | Task |
-Yelp | 31,081 (3) | 411,263 | User-Business | LP |
-Movielens | 2,672 (4) | 234,695 | User-Movie | LP |
-DBLP | 26,128 (4) | 119,783 | Author (4) | NC, CL |
-ACM | 11,246 (3) | 17,426 | Paper (3) | NC, CL |
-IMDB | 11,519 (3) | 17,009 | Movie (3) | NC, CL |
-Freebase | 43,854 (4) | 75,517 | Movie (3) | NC, CL |
-5.1.2. Baselines
-We compare MF2Vec against state-of-the-art heterogeneous graph neural networks (HGNNs), categorized as follows: (a) Relation-Based HGNNs, including RGCN, RGAT, HGT, HetSANN, and Simple-HGN; and (b) Metapath-Based HGNNs, comprising MP2Vec, HAN, MAGNN, HeCo, MHNF, HGMAE HMSG, and MECCH.
-5.1.3. Graph Construction Settings
-Following Algorithm 2, we ensured that the node types matched the target node and that the path length was at most 5.
-For each target node, we performed up to 1000 trials.
-The number of negative samples was set to 1.
-5.1.4. Experimental Settings
-To ensure a fair comparison among all heterogeneous graph neural networks (HGNNs), the node embedding dimension was fixed to 64.
-The learning rate was set to , weight decay to , and the dropout rate to 0.5.
-For each baseline model, hyperparameters such as batch size, attention heads, dropout rate, learning rate, and epochs—excluding the node embedding dimension—followed the default settings specified in the original papers.
-Training was conducted using the Adam optimizer with early stopping (patience of 20 epochs) to prevent overfitting.
-Specifically, for MF2Vec, the temperature () was set to 0.5, and the number of facets () was set to 5.
-To demonstrate generalization, the experiments were conducted using five seed values (1, 10, 100, 1000, and 10000), with the data split into train (80%), validation (10%), and test (10%).
-The model seeds were fixed during the experiments, and the average and standard deviation of the results were calculated for all seeds.
-5.2. Downstream Task (RQ1)
-Dataset | DBLP | ACM | IMDB | Freebase | |||||||||
-Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
-Relation-based | RGCN | 0.8580.004 | 0.8420.005 | 0.8370.005 | 0.8100.012 | 0.7690.007 | 0.7620.011 | 0.5330.024 | 0.4310.027 | 0.3370.043 | 0.5740.049 | 0.5080.057 | 0.4080.078 |
-RGAT | 0.8490.003 | 0.8290.004 | 0.8230.004 | 0.8030.021 | 0.7610.023 | 0.7490.025 | 0.5580.011 | 0.4240.015 | 0.4090.016 | 0.6220.027 | 0.5530.033 | 0.4870.037 | |
-HGT | 0.9260.017 | 0.8890.026 | 0.8990.022 | 0.7790.016 | 0.7410.016 | 0.7090.024 | 0.5400.019 | 0.3960.024 | 0.3860.022 | 0.5780.030 | 0.4820.034 | 0.4270.037 | |
-HetSANN | 0.9500.006 | 0.9310.007 | 0.9250.008 | 0.7800.023 | 0.7360.025 | 0.7110.027 | 0.6130.006 | 0.4870.010 | 0.4840.007 | 0.6460.021 | 0.5750.027 | 0.5210.030 | |
-Simple-HGN | 0.9470.011 | 0.9280.011 | 0.9200.015 | 0.8400.037 | 0.8080.033 | 0.7930.045 | 0.6120.035 | 0.4970.038 | 0.4810.045 | 0.6710.014 | 0.6300.017 | 0.5480.031 | |
-Metapath-based | MP2Vec | 0.5860.020 | 0.3990.025 | 0.3720.026 | 0.7310.010 | 0.6630.010 | 0.6330.014 | 0.5110.011 | 0.3970.016 | 0.3780.013 | 0.6050.019 | 0.6030.022 | 0.5060.030 |
-HAN | 0.9370.003 | 0.9190.003 | 0.9090.004 | 0.7850.011 | 0.7620.009 | 0.7290.014 | 0.5500.020 | 0.4260.021 | 0.3840.034 | 0.6650.011 | 0.6430.014 | 0.5170.033 | |
-MAGNN | 0.9720.003 | 0.9260.003 | 0.9150.003 | 0.9450.014 | 0.8060.031 | 0.7770.049 | 0.7390.013 | 0.5340.032 | 0.4920.070 | 0.7440.015 | 0.6740.019 | 0.5180.063 | |
-HeCo | 0.9780.003 | 0.8940.013 | 0.8860.014 | 0.8310.019 | 0.7450.012 | 0.7030.013 | 0.5440.006 | 0.3850.014 | 0.3600.004 | 0.7530.021 | 0.6520.021 | 0.5150.033 | |
-MHNF | 0.9250.004 | 0.8960.006 | 0.8860.006 | 0.8830.009 | 0.7600.012 | 0.7110.016 | 0.5300.042 | 0.4080.042 | 0.3240.071 | 0.5120.013 | 0.4510.014 | 0.2650.053 | |
-HGMAE | 0.9480.018 | 0.8740.121 | 0.8290.084 | 0.8120.011 | 0.7320.012 | 0.7360.019 | 0.6140.019 | 0.4270.016 | 0.3530.033 | 0.7270.030 | 0.6220.030 | 0.4780.042 | |
-HMSG | 0.9780.003 | 0.9010.020 | 0.8940.021 | 0.9230.013 | 0.7950.019 | 0.7850.020 | 0.6470.021 | 0.4720.017 | 0.4610.022 | 0.7440.026 | 0.6790.040 | 0.5310.038 | |
-MECCH | 0.9910.133 | 0.9410.165 | 0.9370.176 | 0.9160.133 | 0.7900.165 | 0.7930.176 | 0.6570.133 | 0.5120.165 | 0.5020.176 | 0.7730.133 | 0.6480.165 | 0.5880.176 | |
-MF2Vec | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
-5.2.1. Node Classification
-In our node classification experiments, we convert nodes into low-dimensional vector embeddings and apply a Softmax function to obtain probability distributions over label categories.
-We evaluate the model using Micro-F1, Macro-F1, and AUC metrics, providing a balanced assessment of classification performance.
-These metrics are used to compare our model, MF2Vec, against baseline models on the DBLP, ACM, IMDB and Freebase datasets, with results presented in Table LABEL:table_1.
-Our approach consistently outperforms baseline models across all metrics, demonstrating the effectiveness of MF2Vec.
-The significant performance improvement, particularly on the IMDB dataset, highlights the model’s ability to capture complex relational data through multi-facet vectors.
-This suggests that MF2Vec effectively reveals nuanced relationships within heterogeneous information networks (HINs), enhancing understanding beyond node types.
-5.2.2. Link Prediction
-Dataset | Yelp | Movielens | ||||
-Model | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
-RGCN | 0.6350.141 | 0.6210.103 | 0.6140.134 | 0.7410.003 | 0.6640.003 | 0.6640.003 |
-RGAT | 0.6970.104 | 0.6650.081 | 0.6550.086 | 0.6870.115 | 0.6360.101 | 0.6300.104 |
-HGT | 0.8320.004 | 0.7710.004 | 0.7700.005 | 0.7550.136 | 0.7110.116 | 0.7110.115 |
-HetSANN | 0.7030.017 | 0.6890.013 | 0.6880.014 | 0.7740.013 | 0.7720.013 | 0.7710.013 |
-Simple-HGN | 0.8240.016 | 0.7570.018 | 0.7570.018 | 0.8200.007 | 0.7460.012 | 0.7450.015 |
-MP2Vec | 0.7180.003 | 0.6580.002 | 0.6640.003 | 0.7980.004 | 0.6620.004 | 0.6690.002 |
-HAN | 0.6270.102 | 0.5280.076 | 0.3990.140 | 0.8220.004 | 0.6420.050 | 0.6040.072 |
-MAGNN | 0.8350.005 | 0.7730.002 | 0.7740.003 | 0.8580.010 | 0.7760.008 | 0.7810.009 |
-HeCo | 0.8380.003 | 0.7670.004 | 0.7550.002 | 0.8390.003 | 0.7550.004 | 0.7710.005 |
-HMSG | 0.8510.003 | 0.7840.002 | 0.8030.004 | 0.8140.006 | 0.7440.006 | 0.7530.013 |
-MECCH | 0.8670.012 | 0.7820.013 | 0.7740.015 | 0.7520.006 | 0.6900.015 | 0.6340.040 |
-MF2Vec | 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 |
-In Movielens and Yelp, link prediction is framed as a binary classification task to determine the presence of an edge in the original graph.
-During training, validation, and testing, negative edges are created by replacing the destination node in positive edges with a randomly chosen node that is not connected to the source node.
-The link probability is computed using the dot product of node embeddings: , where is the sigmoid function.
-Table 4 presents the AUC, Micro-F1, and Macro-F1 scores for our model and baselines on Yelp and Movielens.
-MF2Vec performs competitively on Yelp (against MECCH) and outperforms all baselines on MovieLens.
-These results highlight the advantage of using multi-facet vectors derived from node paths for link prediction, and demonstrate the efficacy of our model in updating embeddings without relying on node types.
-Dataset | DBLP | ACM | IMDB | Freebase | ||||
-Model | NMI | ARI | NMI | ARI | NMI | ARI | NMI | ARI |
-RGCN | 0.0168 | 0.0670 | 0.3612 | 0.2809 | 0.0088 | 0.0056 | 0.0228 | 0.0443 |
-RGAT | 0.4435 | 0.4097 | 0.3976 | 0.3140 | 0.0477 | 0.0143 | 0.0327 | 0.0387 |
-HGT | 0.7722 | 0.8250 | 0.5064 | 0.5335 | 0.0433 | 0.0459 | 0.0801 | 0.1075 |
-HetSANN | 0.752 | 0.7723 | 0.5274 | 0.5263 | 0.0865 | 0.043 | 0.1054 | 0.0967 |
-Simple-HGN | 0.7707 | 0.6405 | 0.5309 | 0.5237 | 0.1410 | 0.1180 | 0.1541 | 0.1804 |
-MP2Vec | 0.0519 | 0.0330 | 0.1034 | 0.0641 | 0.0065 | 0.0096 | 0.0145 | 0.0286 |
-HAN | 0.3533 | 0.1845 | 0.3036 | 0.3528 | 0.0057 | 0.0015 | 0.0211 | 0.0217 |
-MAGNN | 0.7964 | 0.8193 | 0.5125 | 0.5317 | 0.1075 | 0.1199 | 0.1872 | 0.1477 |
-HeCo | 0.7477 | 0.7872 | 0.5394 | 0.5454 | 0.0049 | 0.0028 | 0.0805 | 0.0793 |
-MHNF | 0.7197 | 0.7498 | 0.5385 | 0.5274 | 0.1443 | 0.1467 | 0.1881 | 0.1697 |
-HMSG | 0.7278 | 0.7767 | 0.4475 | 0.3630 | 0.0335 | 0.0326 | 0.1257 | 0.1213 |
-MECCH | 0.8312 | 0.8739 | 0.4202 | 0.3513 | 0.0070 | 0.0056 | 0.2612 | 0.3127 |
-MF2Vec | 0.8606 | 0.8951 | 0.5714 | 0.5790 | 0.1459 | 0.1448 | 0.3083 | 0.3729 |
-5.2.3. Node Clustering
-We carry out node clustering experiments using the DBLP, ACM, IMDB, and Freebase datasets to assess the embedding quality produced by MF2Vec.
-In these experiments, similar to node classification, we utilize the embedding vectors of target nodes from the testing set as input for the K-Means model and employ NMI and ARI metrics to gauge performance.
-As presented in Table 5, our model surpasses the baseline models, except for MHNF (Sun et al., 2022) in ARI on the IMDB dataset, where MF2Vec closely follows it by a small margin, demonstrating our model’s superior ability to generate effective node representations based on multi-facet vectors for node clustering.
-5.3. Stability of performance (RQ2)
-Dataset | Metric | MAGNN | HMSG | MECCH | MF2Vec |
-DBLP | AUC | 0.144 | 0.201 | 0.118 | 0.064 |
-Mi-F1 | 0.233 | 0.318 | 0.195 | 0.138 | |
-Ma-F1 | 0.251 | 0.277 | 0.197 | 0.139 | |
-ACM | AUC | 0.147 | 0.188 | 0.211 | 0.141 |
-Mi-F1 | 0.083 | 0.106 | 0.076 | 0.070 | |
-Ma-F1 | 0.102 | 0.088 | 0.066 | 0.051 | |
-IMDB | AUC | 0.042 | 0.041 | 0.043 | 0.031 |
-Mi-F1 | 0.029 | 0.049 | 0.067 | 0.028 | |
-Ma-F1 | 0.032 | 0.030 | 0.032 | 0.020 |
-This section emphasizes the stability advantage of leveraging multi-facets from paths in our model.
-Unlike metapath-based models that depend on predefined schemas (e.g., APA, APCPA), MF2Vec achieves consistent performance across different schema types, effectively capturing fine-grained information regardless of metapath type.
-To validate this, we compare single-metapath performance across all models, including MF2Vec, on the DBLP, ACM, and IMDB datasets.
-Table 6 presents the standard deviations in node classification performance using a single metapath.
-While other models show significant performance variation depending on the chosen metapath, MF2Vec demonstrates remarkable stability.
-For the DBLP dataset, performance variations highlight the reliance of metapath-based models on predefined paths and the associated time and cost of analysis, which MF2Vec effectively mitigates.
-5.4. Time Complexity (RQ3)
-Model | Time Complexity |
-HetSANN | |
-SimpleHGN | |
-MAGNN | |
-HAN | |
-MECCH | |
-MF2Vec |
-The proposed model’s time complexity is , where is the number of nodes, is the average neighbors, is the number of facets, and is the embedding dimension.
-Attention adds , and graph convolution contributes , resulting in a total complexity of .
-In Table 7, models like MECCH (Fu and King, 2024), HAN (Wang et al., 2019), and MAGNN (Fu et al., 2020) rely on metapaths, incurring higher costs due to growth with metapath length ().
-MF2Vec avoids this by leveraging facets (), leading to more efficient computation, satisfying .Additionally, we compare MF2Vec with relation-based methods like HetSANN (Hong et al., 2020) and SimpleHGN (Lv et al., 2021), which also avoid metapaths by focusing on edge types.
-Figure 5 compares training times (log scale) and Macro F1 scores for MF2Vec and other models on DBLP, IMDB, and Movielens.
-MAGNN (Fu et al., 2020) and MP2Vec (Dong et al., 2017) are excluded from Movielens due to their excessive training times.
-MF2Vec achieves the best performance on node classification and link prediction tasks, with significantly lower training times.
-This efficiency arises from its dynamic use of path characteristics, avoiding the high costs of aggregating multiple paths as seen in MAGNN and MP2Vec, ensuring better scalability and faster convergence without compromising performance.
-5.5. Visualization
-To evaluate embedding quality, we visualize node representations from various HGNNs on the DBLP dataset using t-SNE (Van der Maaten and Hinton, 2008).
-Figure 6 shows t-SNE plots for HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MECCH (Fu and King, 2024), and MF2Vec, with colors representing author classes.
-MF2Vec demonstrates superior clustering within classes and clear decision boundaries, outperforming the more scattered intra-class distributions and weaker inter-class separations of other HGNNs.
-6. Study of MF2Vec
-To analyze the impact of intrinsic characteristics of MF2Vec with respect to various features including train rati0, the number facets, we additionally conducted extensive experiments.
-Dataset | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
-Method | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 |
-No.
-| 0.8130.003 | 0.7430.002 | 0.7520.004 | 0.8320.003 | 0.7540.005 | 0.7430.013 | 0.9740.004 | 0.8680.012 | 0.8540.012 | 0.9120.014 | 0.7880.015 | 0.7590.025 | 0.531 0.022 | 0.378 0.005 | 0.232 0.036 | 0.7310.011 | 0.6460.021 | 0.4750.015 |
-Random.
-| 0.8670.018 | 0.7970.003 | 0.8080.015 | 0.8710.006 | 0.7950.006 | 0.8030.005 | 0.9890.003 | 0.9350.009 | 0.9310.01 | 0.9490.009 | 0.8240.013 | 0.8130.012 | 0.7510.005 | 0.5880.004 | 0.5710.003 | 0.8250.020 | 0.6960.017 | 0.6280.042 |
-Gumble.
-| 0.8930.003 | 0.8200.005 | 0.8240.004 | 0.9180.003 | 0.8400.004 | 0.8470.007 | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
-6.1. Train Ratio ()
-As shown in Figure 7, MF2Vec consistently achieves the highest performance across varying training set ratios, demonstrating its robustness and adaptability.
-While the performance of all models declines as the training ratio decreases, MF2Vec remains its superiority in most cases and retains competitive in the rest.
-At the lowest ratio of 0.2, it occasionally shows comparable performance to other models.
-HAN and one baseline were excluded due to significantly lower performance, ensuring clearer visualization.
-These results highlight MF2Vec’s effectiveness in handling sparse training data.
-6.2. Number of Facets ()
-Figure 8 shows MF2Vec’s performance across six datasets with varying facet counts.
-The performance difference relative to “” highlights the advantage of multiple facets.
-Except for IMDB, optimal results are achieved with five facets, consistently outperforming the single-facet case by effectively leveraging node and path information.
-Beyond five facets, performance declines, likely due to reduced feature grouping efficiency.
-6.3. Role of Facet Selection
-As MF2Vec selects the optimal facet for each path in a learnable manner, we conduct an additional experiment to validate its facet selection strategy.
-Table 8 compares different methods for selecting facet information as edge weights.
-Our GCN-based results show that the proposed method, using Gumbel-Softmax and attention mechanisms, outperforms all alternatives across datasets.
-Random facet selection reduces performance, and omitting facets entirely leads to the lowest results.
-These findings highlight the importance of facet-related information for effective graph learning and the proposed method’s ability to capture complex structures and hidden relationships.
-7. Case Study
-To evaluate facet balance among node types, we analyzed their distribution as shown in Figure 9.
-The heatmaps reveal equitable attention weights across node types, with each column representing a specific node and highlighting diverse facet distributions.
-For example, the Paper node “P1612” emphasizes Facet 1, while the Author node “A3958” distributes attention evenly across facets, reflecting varying significance assigned to facets based on connections.
-This analysis confirms that nodes either focus on specific facets or distribute attention uniformly, with no strong bias toward any facet, validating the importance of modeling shared facets in heterogeneous graphs.
 6. Conclusion
-In this study, we introduce MF2Vec, a novel heterogeneous graph neural network model, which generates node representations using multi-facet vectors from various paths, regardless of node types.
-MF2Vec captures complex latent semantics through these multi-faceted paths, offering a node-type-agnostic embedding approach.
-Evaluations across six datasets demonstrate that MF2Vec outperforms leading baseline models in both performance and time complexity.
-Future work could focus on automating path selection within heterogeneous graph neural networks using learnable methods to improve upon current random pre-selection methods, highlighting the potential of dynamic multi-facet path utilization in graph networks.
-Appendix
-A. Baselines Code and Datasets
-MF2Vec is implemented using PyTorch. Table 9 lists the URLs of the authors’ implementations of the compared methods, and Table 10 provides the URLs where the datasets can be downloaded.
-B. Baselines
-We compare MF2Vec against state-of-the-art relation-based HGNNs and metapath-based HGNNs , categorized as follows:
-Relation-Based HGNNs.
-RGCN: Introduces a relational graph convolutional network that assigns weights to each relation type, enabling the learning of structural information in heterogeneous graphs.
-RGAT: Applies attention mechanisms to each relation type in relational graphs, effectively integrating information in heterogeneous graphs.
-HGT: Utilizes a modified transformer architecture that considers various node and edge types in heterogeneous graphs to learn complex relationships.
-HetSANN: Applies attention mechanisms to learn structural information between nodes in heterogeneous graphs without relying on metapaths.
-Simple-HGN: Provides efficient and effective node embeddings in heterogeneous graphs through a simplified structure.
-Metapath-Based HGNNs.
-MP2Vec: Proposes a node embedding technique in heterogeneous networks using metapaths to effectively capture diverse relationships.
-HAN: Develops a hierarchical attention mechanism-based model to learn the importance of nodes in heterogeneous graphs using metapaths.
-MAGNN: Enhances node embedding performance in heterogeneous graphs by leveraging intermediate node information within metapaths.
-HeCo: Combines self-supervised learning and contrastive learning in heterogeneous graphs to improve node representation learning performance.
-MHNF: A framework that automatically extracts metapaths in heterogeneous networks and utilizes them for node embedding learning.
-HGMAE: Utilizes masked autoencoders to learn robust node representations in heterogeneous graphs.
-HMSG: Integrates multi-scale structural information in heterogeneous graphs to enhance node representation learning performance.
-MECCH: Improves node embedding performance in heterogeneous graphs by leveraging contextual information of metapaths.
-C. Stability of Performance Across Metapaths
-Tables 11, 12, and 13 present performance metrics for the top three models and MF2Vec across the DBLP, ACM, and IMDB datasets, respectively, based on a single metapath.
-Across all datasets, MF2Vec consistently outperforms other models, achieving the highest Mi-F1, Ma-F1, and AUC scores for all metapaths.
-On the DBLP dataset (Table 11), MF2Vec excels across all metrics.
-Similarly, on the ACM dataset (Table 12), it achieves the best performance for both PAP and PSP metapaths.
-On the IMDB dataset (Table 13), MF2Vec again demonstrates its robustness and consistency, achieving the highest scores across all metapaths.
-DBLP | ||||||||||||
-Metapath | APTPA | APCPA | APA | ALL | ||||||||
-Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
-MAGNN | 0.670 | 0.685 | 0.912 | 0.918 | 0.925 | 0.988 | 0.387 | 0.434 | 0.683 | 0.910 | 0.921 | 0.987 |
-HMSG | 0.604 | 0.498 | 0.843 | 0.889 | 0.887 | 0.990 | 0.345 | 0.265 | 0.563 | 0.929 | 0.925 | 0.988 |
-MECCH | 0.825 | 0.819 | 0.946 | 0.936 | 0.933 | 0.993 | 0.527 | 0.529 | 0.746 | 0.953 | 0.951 | 0.994 |
-MF2Vec | 0.904 | 0.907 | 0.986 | 0.958 | 0.956 | 0.990 | 0.667 | 0.670 | 0.861 | 0.963 | 0.962 | 0.995 |
-ACM | |||||||||
-Metapath | PAP | PSP | ALL | ||||||
-Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
-MAGNN | 0.6064 | 0.6602 | 0.8754 | 0.7298 | 0.7649 | 0.8575 | 0.8095 | 0.8254 | 0.9480 |
-HMSG | 0.6608 | 0.6024 | 0.8598 | 0.7656 | 0.7128 | 0.8302 | 0.8204 | 0.8160 | 0.9151 |
-MECCH | 0.6683 | 0.6405 | 0.8431 | 0.7531 | 0.7067 | 0.8440 | 0.7980 | 0.7922 | 0.9194 |
-MF2Vec | 0.8080 | 0.8013 | 0.9322 | 0.7631 | 0.7211 | 0.8436 | 0.8653 | 0.8610 | 0.9528 |
-IMDB | |||||||||
-Metapath | MAM | MDM | ALL | ||||||
-Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
-MAGNN | 0.5241 | 0.5276 | 0.7072 | 0.5057 | 0.5180 | 0.6971 | 0.5685 | 0.5731 | 0.7493 |
-HMSG | 0.4717 | 0.4587 | 0.6213 | 0.4226 | 0.3634 | 0.5554 | 0.4767 | 0.4357 | 0.6306 |
-MECCH | 0.4149 | 0.3169 | 0.5887 | 0.4029 | 0.3601 | 0.5603 | 0.4639 | 0.4404 | 0.6449 |
-MF2Vec | 0.5875 | 0.5733 | 0.7180 | 0.5851 | 0.5565 | 0.7142 | 0.6211 | 0.6117 | 0.7709 |
-Model | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
-AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
-GCN (Ours) | 0.8947 | 0.8200 | 0.8261 | 0.9074 | 0.8283 | 0.8366 | 0.9958 | 0.9631 | 0.9623 | 0.9528 | 0.8653 | 0.8610 | 0.7709 | 0.6211 | 0.6117 | 0.8701 | 0.7429 | 0.6991 |
-Edge GAT | 0.8624 | 0.8041 | 0.7941 | 0.8629 | 0.8086 | 0.8180 | 0.9900.002 | 0.9280.007 | 0.9210.007 | 0.9500.005 | 0.8540.008 | 0.8490.010 | 0.7450.013 | 0.5790.023 | 0.5580.025 | 0.7950.010 | 0.6810.010 | 0.6140.026 |
-GraphSAGE | 0.8296 | 0.7701 | 0.7771 | 0.8028 | 0.7506 | 0.7719 | 0.9830.005 | 0.9020.020 | 0.8960.022 | 0.9380.006 | 0.8160.009 | 0.8090.009 | 0.7020.018 | 0.5310.029 | 0.5010.031 | 0.7340.018 | 0.6400.021 | 0.4710.016 |
-We evaluate the performance of these models across three downstream tasks: node classification, link prediction, and node clustering.
-D. Ablation Study
-D.1. Effect of Temperature ()
-From Figure 10, we observe the sensitivity of the Gumbel-Softmax in calculating the probability across all classes with varying .
-According to the performance metrics, the model’s performance exhibits minor variations in Movielens, DBLP, and Yelp, and changes within a 2% range in Movielens and Freebase.
-The results indicate that the sensitivity of the model to different values of is relatively stable across datasets, with minor fluctuations in the Macro F1 and Micro F1 scores.
-D.2. Graph Method
-Table 14 compares GCN, EdgeGAT, and GraphSAGE to evaluate the impact of facet features as edge weights.
-GCN outperforms across all datasets, achieving AUC and F1 scores above 0.99 on DBLP and ACM by effectively leveraging edge weights to capture local and global relationships.
-While EdgeGAT models relationship importance using attention, its performance is less consistent, and GraphSAGE shows lower results due to its limited capacity to represent multi-facet relationships despite its efficiency.
-These findings highlight the significance of edge weights in facet-based graph learning.
-F. Algorithms for Random Walk and Sub-graph Generation
-This section describes the algorithms used for random walk and sub-graph generation in heterogeneous networks.
-The random walk algorithm explores paths within the graph, while the sub-graph generation algorithm utilizes these paths to construct sub-graphs based on specific node types.
-F.1. Random Walk Algorithm
-The first algorithm 2 performs a random walk on a network.
-It generates a path of a specified length starting from a given node.
-The path is built by selecting a random neighbor at each step.
-If a neighbor cannot be found or if the end node of the path does not match the desired node type, the algorithm returns None.
-F.2. Sub-graph Generation Using Random Walk Paths
-The second algorithm 3 generates a sub-graph from the network using multiple random walks.
-For each node, several random walks are performed to collect paths.
-Unique node pairs are recorded to avoid duplicates.
-The sub-graph is constructed by adding nodes and edges based on the collected paths.
+We introduced MF2Vec, a new heterogeneous graph neural network that learns node representations using multi-facet paths, independent of node types.
+MF2Vec consistently outperforms existing methods in node classification, link prediction, and clustering across six benchmark datasets.
+It demonstrates high stability against meta-path variations and offers improved computational efficiency.
+Our ablation results confirm the effectiveness of learnable facet selection.
+In future work, we plan to automate facet and path selection to further improve adaptability and performance.
+While MF2Vec shows strong generalization and scalability, its current path selection is based on random walks, which may miss some informative paths.
+Addressing this limitation is an important direction for future research.
+MF2Vec offers a simple and scalable solution for heterogeneous graph learning.
+Acknowledgements.
+This research was supported by National Research Foundation of Korea (NRF-2022M3J6A1063021).GenAI Usage Disclosure
+The authors used Generative AI tools (ChatGPT, OpenAI) only for minor grammar checking and assistance in generating table formatting.
+No part of the scientific content, analysis, or results was produced by GenAI.